### v1

# Homogeneous maximizers of the Blaschke–Santaló-type functionals

Abstract.
We study Blaschke–Santaló-type inequalities for $N\geq 2$ sets (functions) and a special class of cost functions. In particular, we prove new results about reduction of the maximization problem for the Blaschke–Santaló-type functional to homogeneous case (functional inequalities on the sphere) and extend the symmetrization argument to the case of $N>2$ sets. We also discuss links to the multimagrinal optimal transportation problem and the related sharp transportation-information inequalities.
1. Introduction
In this work we discuss various extensions of the classical Blaschke–Santaló inequality
(1) $|A||A^{\circ}|\leq|B_{2}|^{2}$
and its functional counterpart
(2) $\int_{\mathbb{R}^{n}}e^{-V}dx\int_{\mathbb{R}^{n}}e^{-V^{*}}dy\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{|x|^{2}}{2}}dx\Bigr)^{2}.$
Here $A\subset\mathbb{R}^{n}$ is a convex symmetric (A=A-A=A) body and $V\colon\mathbb{R}^{n}\to(-\infty,\infty]$ is a proper convex even function,
$A^{\circ}=\{y:\langle x,y\rangle\leq 1\}$
is the polar set of $A$ and
$V^{*}(y)=\sup_{x}(\langle x,y\rangle-V(x))$
is the Legendre transform of $V$. Finally, $B_{p}=\{x:\sum_{j=1}^{n}|x_{j}|^{p}\leq 1\}\subset\mathbb{R}^{n}$ is the $l^{p}$-unit ball and $|\cdot|$ is the Lebesgue volume. Inequality (1) was proved by Santaló [36] (and by Blaschke in dimension 3). Inequality (2) is substantially younger and was proved first by K. Ball [4] in 1986.
Let us consider $N\geq 2$ copies of $\mathbb{R}^{n}$. The $i$-th copy will be equipped with the standard coordinate system $x_{i}$. We write
$x=(x_{1},\cdots,x_{N})\in(\mathbb{R}^{n})^{N},\ x_{i}\in\mathbb{R}^{n}.$
The coordinate representation of arbitrary $x_{i}\in\mathbb{R}^{n}$ will be written in the following way:
$x_{i}=(x_{i,1},\cdots,x_{i,j},\cdots,x_{i,n}).$
We prove the following generalization of the both versions of the Blaschke–Santaló inequality.
Theorem 1.1.
Let $N\geq 2$ be a natural number. Consider the following cost function
$c(x)=c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$
on $(\mathbb{R}^{n})^{N}$. Assume we are given $N$ symmetric convex sets $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ such that $c(x_{1},\cdots,x_{N})\leq 1$ on the set $\prod_{i=1}^{N}A_{i}\subset(\mathbb{R}^{n})^{N}$. Then
$\prod_{i=1}^{N}|A_{i}|\leq|B_{N}|^{N}.$
Let $\{V_{i}\}$, $1\leq i\leq N,$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$. Then
$\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}dy\Bigr)^{N}.$
In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
This example can be considered as an illustration of several results of this work, where the central question is the relation between the functional and the set version of the Blaschke–Santaló type functional. We can put it into the following form: when the Blaschke–Santaló functional admits homogeneous extremizers for given homogeneous reference measures and homogeneous cost function?
Before we explain the other results of this paper, let us give an incomplete overview of the related previous results below. See more information in the nice recent survey of Fradelizi, Meyer and Zvavitch [20].
  • Generalizations to the non-symmetric case, proofs of the classical case.
    It was proved in Artstein–Klartag–Milman [1] that for any convex (not necessary symmetric) $V$ there exists a point $a\in\mathbb{R}^{n}$ such that
    (3) $\int_{\mathbb{R}^{n}}e^{-\tilde{V}}dx\int_{\mathbb{R}^{n}}e^{-\tilde{V}^{*}}dy\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{|x|^{2}}{2}}dx\Bigr)^{2},$
    where $\tilde{V}(x)=V(x-a)$. If $\int_{\mathbb{R}^{n}}xe^{-V}dx=0$, one can take $a=0$.
    Let us give a short comment about the proofs in the classical case. A short proof of (2) not relying on the set version (1) was obtained by Lehec [28]. Bianchi and Kelli [6] proved the classical Blaschke–Santaló inequality using Fourier analysis. Nakamura and Tsuji [31] obtained a proof of the functional Blaschke–Santaló inequality based on the heat flow approach (see a non-symmetric generalization in Cordero-Erausquin–Fradelizi–Langharst [13] and a more direct proof in Cordero-Erausquin–Gozlan–Nakamura–Tsuji [14]). Moreover, they obtained a more general functional $L^{p}-L^{q}$ inequality. A variational proof of (2) based on the uniqueness of solution to the corresponding Monge–Ampère equation has been obtained in Colesanti–Kolesnikov–Livschyts–Rotem [11] (see also subsection 3.1 here). For questions of stability see Barthe–Böröczky–Fradelizi[5] and the reference therein.
  • Generalized Legendre-type transforms.
    Some other type of dualities for functions and related Blaschke–Santaló-type inequalities have been considered already in [1]. An important extension, which is particularly relevant to our work is the result of Fradelizi and Meyer [19]. They established inequality of the type
    $\int_{\mathbb{R}^{n}}f(x)dx\int_{\mathbb{R}^{n}}g(y)dy\leq\Bigl(\int_{\mathbb{R}^{n}}\rho(|x|^{2})dx\Bigr)^{2}$
    under the constraint $f(x)g(y)\leq\rho(\langle x-z,y-z\rangle)$, where $f,g\colon\mathbb{R}^{n}\to\mathbb{R}_{+}$ and $\rho\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ (see [19] for precise formulation). Another result of [19] is the following generalization of (2):
    (4) $\int_{\mathbb{R}^{n}}\rho(V)dx\int_{\mathbb{R}^{n}}\rho(V^{*})dy\leq\Bigl(\int_{\mathbb{R}^{n}}\rho\Bigl(\frac{|x|^{2}}{2}\Bigr)dx\Bigr)^{2}.$
    Here $\rho$ is non-increasing and log-concave, $V$ is even and convex.
    There are several extensions of the Blaschke–Santaló inequality for special types of the generalized Legendre transform. See, for instance, Rotem [35], Artstein–Slomka [3], Artstein–Sadowski–Wyczesany [2], Fradelizi–Gozlan–Sadovsky–Zugmeyer [18].
  • Relation to optimal transportation and information theory.
    It is well–known that inequality (2) admits a transportation counterpart. This can be established by the arguments going back to B. Maurey (see M. Fathi [17]): let $\gamma$ be the standard Gaussian measure and $f\cdot\gamma,g\cdot\gamma$ be probability measures such that $f\cdot\gamma$ is centered: $\int_{\mathbb{R}^{n}}xfd\gamma=0$. Then the following transportation inequality holds
    (5) $\frac{1}{2}W^{2}_{2}(f\cdot\gamma,g\cdot\gamma)\leq\int_{\mathbb{R}^{n}}f\log fd\gamma+\int_{\mathbb{R}^{n}}g\log gd\gamma.$
    Here $W_{2}$ is the transportation distance on probability measures
    $W^{2}_{2}(\mu_{1},\mu_{2})=\inf_{\pi\in\Pi(\mu_{1},\mu_{2})}\int_{(\mathbb{R}^{n})^{2}}|x-y|^{2}\pi(dxdy),$
    where $\Pi(\mu_{1},\mu_{2})$ is the set of probability measures with projections $\mu_{1},\mu_{2}$. Inequality (5) can be extracted from (2) and vice versa by relatively simple arguments. Later Gozlan [22] extended the arguments of Fathi and established a transportation version of the famous Mahler conjecture. Following observation of Kolesnikov [25] about transportation functionals on the sphere Fradelizi–Gozlan–Sadovsky–Zugmeyer [18] proved a variant of inequality (5) on $\mathbb{S}^{n-1}$ for appropriate cost function (see also [7]). Courtade–Fathi–Mikulincer obtained in [15] a stochastic proof of (5).
  • Generalization for $N>2$ functions.
    Motivated by the theory of geodesic barycenters (”Wasserstein barycenters”) of probability measures, Kolesnikov and Werner [27] conjectured that the following inequality might be true:
    (6) $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{|x|^{2}}{2}}dx\Bigr)^{N}$
    under assumption that functions $\{V_{i}(x_{i})\}$ are even and satisfy
    $\sum_{i=1}^{N}V_{i}(x_{i})\geq\frac{1}{N-1}\sum_{j\neq i}\langle x_{i},x_{j}\rangle.$
    The choice of the cost function
    $c_{bar}=\sum_{j\neq i}\langle x_{i},x_{j}\rangle$
    is motivated by the barycenter problem. Clearly, (6) is a generalization of (2).
    Kolesnikov and Werner [27] proved that (6) holds under assumption of unconditionality. Some generalizations of this result were obtained by Kalantzopoulos and Saroglou [23]. In particular, they conjectured the following set version of (6):
    (7) $\prod_{i=1}^{N}|A_{i}|\leq|B_{2}|^{N},$
    where $A_{i}$ are convex symmetric sets satisfying $\sum_{j\neq i}\langle x_{i},x_{j}\rangle\leq\frac{N(N-1)}{2}$ for any tuple $\{x_{i}\}$ such that $x_{i}\in A_{i}$. The question about inequality (7) remains open to our best knowledge.
    Inequality (6) was proved in full generality by Nakamura and Tsuji [32]. Finally, Courtade and Wang [16] obtained a generalization of (6) in the non-symmetric case.
  • Weighted Blaschke–Santaló inequality.
    Another generalization of (2) was obtained in Colesanti–Kolesnikov–Livschyts–Rotem [11]. They studied inequalities of the following type
    (8) $\Bigl(\int_{\mathbb{R}^{n}}e^{-V}dx\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{p-1}V^{*}(\nabla\Phi)}dx\Bigr)^{\frac{p-1}{p}}\leq\int_{\mathbb{R}^{n}}e^{-\Phi}dx,$
    where $\Phi$ is a $p$-homogeneous convex even function and $p>1$.
    Assume we are given a strictly convex (and sufficiently regular) $p$-homogeneous even function $\Phi$ such that (8) holds for every even $V$. $p$-homogeneity implies that $\Phi=\frac{1}{1-p}\Phi^{*}(\nabla\Phi)$, hence $\Phi$ is a maximum point of the functional
    $\mathcal{F}_{p,\Phi}(V)$ $=\Bigl(\int_{\mathbb{R}^{n}}e^{-V}dx\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{p-1}V^{*}(\nabla\Phi)}dx\Bigr)^{\frac{1}{q}}$
    $=\Bigl(\int_{\mathbb{R}^{n}}e^{-V}dx\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{p-1}V^{*}}\det D^{2}\Phi^{*}dy\Bigr)^{\frac{1}{q}}.$
    We observe that in general the maximizer $\Phi$ corresponds to the non-Gaussian distribution $\frac{e^{-\Phi}}{\int e^{-\Phi}dx}dx$.
    It was realized in [11] that inequality (8) can indeed hold for some functions $\Phi$. The simplest (non-Gaussian) example is given by
    $\Phi=\sum_{i=1}^{n}|x_{i}|^{p}$
    with $p>2$. On the other hand, the same function with $p<2$ does not satisfy (8).
  • Relation to other problems.
    Finally, we mention several interesting developments related to inequalities of the Blaschke–Santaló type. First, motivated by the Nakamura–Tsuji result [32], E. Milman [30] obtained a new proof of the Gaussian correlation inequality, which was a long-standing conjecture. Nakamura and Tsuji [33] generalized his proof and obtained a non-symmetric version of the Gaussian correlation inequality. Another interesting development concerns relation of the Mahler conjecture to the Bourgain slicing problem, see Klartag [24], Fradelizi–Sola [21], and Mastrantonis–Rubinstein [29]. Kolesnikov–Livschyts–Rotem studied in [26] infinitesimal versions of (8) in relation to the log-Brunn–Minkowski problem.
Motivated by these results, we study the following generalized Blaschke–Santaló functional:
(9) $\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})=\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}m_{i}(dx_{i})\Bigr)^{\frac{1}{\alpha_{i}}}.$
Here $m_{i}$ are reference measures on $\mathbb{R}^{n}$ and $\alpha_{i}$ are given positive numbers.
As in the classical case we always assume that functions $V_{i}$ satisfy inequality
$\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x_{1},\cdots,x_{N})=c(x)$
for a given cost function $c$.
In this work we study, in particular, the following questions.
  • A)
    Equivalence between the set and the functional version.
    The equivalence of the set inequality (1) and the functional inequality (2) is well known. Implication (2) $\Longrightarrow$ (1) is almost immediate (one takes $V=\frac{1}{2}|x|^{2}_{K},V^{*}=\frac{1}{2}|x|^{2}_{K^{\circ}}$, where $|\cdot|_{K}$ is the Minkowski functional of $K$ and apply lawyer-cake decomposition). Implication (1) $\Longrightarrow$ (2) is a bit more tricky, the reader can find it, for instance, in [1], [4].
    For the case $N>2$ and the barycentric cost function $c_{bar}$ the equivalence between (6) and (7) is not known.
  • B)
    Gaussianity and homogeneity of maximizers
    A remarkable observation of Nakamura and Tsuji was the fact that inequality (6) fits into a large family of inequalities possessing Gaussian maximizers (the so-called Gaussian saturation principle). Inequalities of this type are called Brascamp–Lieb inequalities and they go back to the classical work of Brascamp and Lieb [9].
    A relevant observation from Colesanti–Kolesnikov–Livschyts–Rotem [11] was homogeneity of the maximizers of $\mathcal{F}_{p,\Phi}$. This property has certain correspondence to the Gaussian saturation principle for inequalities of the Brascamb–Lieb type, but seems to be even more general, because can potentially be applied to situations, where Gaussian maximizers are not possible. An example is given by inequality (8) for $\Phi=\sum_{i=1}^{n}|x_{i}|^{p},p>2$, proved in [11].
  • C)
    Optimal transportation counterpart for the Blaschke–Santaló-type inequalities for $\mathcal{F}_{p,\Phi}$.
    We are interested in generalizations of the sharp Talagrand inequality (5) for the case of $N>2$ functions and non-quadratic costs. Some developments in this direction have been obtained in [17], [27], [22], [18], [15]. This part is strongly related to the multimarginal and barycentric transportation problems, in particular, to duality theory and multiple Legendre-type transforms.
We want to find sufficient conditions ensuring that maximizers of (9) are homogeneous. First we assume that the reference measures $m_{i}=\rho_{i}(x_{i})dx_{i},\ 1\leq i\leq N,$ admit $r_{i}$-homogeneous densities
(10) $\rho_{i}(tx_{i})=t^{r_{i}}\rho_{i}(x_{i}),$
where $r_{i}$ are some nonnegative numbers.
Of course, another natural assumption is that the cost function $c$ is homogeneous as well. The minimal requirement for $c$ will be $p$-homogeneity with respect to $x=(x_{1},\cdots,x_{N})$:
(11) $c(tx)=t^{p}c(x)$
for some positive number $p>0$.
Assume that the maximizers $\{V_{i}(x_{i})\}$ of (9) are $\beta_{i}$-homogeneous
(12) $V_{i}(tx_{i})=t^{\beta_{i}}V_{i}(x_{i}),\ \ \ t\geq 0.$
First we observe a natural consistency condition for numbers $\alpha_{i},\beta_{i},r_{i},p$. We observe that $\beta_{i}$-homogeneity of $V_{i}$ and $p$-homogeneity of $c$ immediately implies that for every $t>0$ one has
$\sum_{i=1}^{N}V_{i}(t^{1-\frac{p}{\beta_{i}}}x_{i})\geq c(x),$
thus the tuple $\{V_{i,t}\}$, where $V_{i,t}(x_{i})=V_{i}(t^{1-\frac{p}{\beta_{i}}}x_{i})$ is admissible. Next we compute using $r_{i}$-homogeneity of $\rho_{i}$ and change of variables:
$\mathcal{BS}_{\alpha,m}(V_{1,t},\cdots,V_{N,t})=t^{\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}\bigl(\frac{p}{\beta_{i}}-1\bigr)}\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N}).$
From this relation we see that $\{V_{i}\}$ can not be a maximizer unless $\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}\Bigl(\frac{p}{\beta_{i}}-1\Bigr)=0.$ Thus it is necessary to assume the following consistency condition :
(13) $p=\frac{\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}}{\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}\beta_{i}}}.$
We note that (13) is a minimal requirement for existence of maximizers, but it is not sufficient (see Example 5.15 in [11]). Another observation is that for many homogeneous cost functions there exist a limited choice of $\beta_{i}$ such that the set of admissible $\beta_{i}$-homogeneous tuples $\{V_{i}\}$ is not empty. Consider, for instance, the barycentric function $c_{bar}(x)=\frac{1}{N-1}\sum_{i,j=1}^{N}\langle x_{i},x_{j}\rangle.$ The reader can easily verify that in this case inequality $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x)$ is possible for $\beta_{i}$-homogeneous $V_{i}$ only if $\beta_{i}=2$.
This is why in addition to (13) we will assume another consistency condition:
(14) $\frac{n+r_{1}}{\beta_{1}}=\frac{n+r_{2}}{\beta_{2}}=\cdots=\frac{n+r_{N}}{\beta_{N}}.$
Under (14) condition (13) can be equivalently rewritten in the following way:
(15) $p=\frac{\sum_{i=1}^{N}\frac{\beta_{i}}{\alpha_{i}}}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}.$
In this work, we extend several results regarding the Blaschke–Santaló inequality in different forms to the following class of cost functions.
Definition 1.2.
We say that a function $c(x_{1},\cdots,x_{N})$ is $(p_{1},\cdots,p_{N})$-homogeneous if for all $1\leq i\leq N$ and $t\geq 0$ one has
$c(x_{1},\cdots,tx_{i},\cdots,x_{N})=t^{p_{i}}c(x).$
Another central result of this work is the following Theorem about homogeneity of maximizers of the Blaschke–Santaló functional.
Theorem 1.3.
Let $m_{i}$ satisfy assumption (10) and have densities $\rho_{i}$ that are positive almost everywhere. Let $c$ be a cost function with the following properties:
  • $c$ is a continuous $(p_{1},\cdots,p_{N})$-homogeneous for some numbers $p_{i}>0$
  • for every $1\leq i\leq N$ there exist numbers $s_{j}\in\{-1,1\}$, $j\neq i$ such that
    $c(x_{1},\cdots,x_{i-1},-x_{i},x_{i+1},\cdots,x_{N})=c(s_{1}x_{1},\cdots,s_{i-1}x_{i-1},x_{i},s_{i+1}x_{i+1},\cdots,s_{N}x_{N})$
  • $c$ is convex with respect to every variable, i.e.
    $x\to c(x_{1},\cdots,x_{i-1},x,x_{i+1},x_{N})$
    is a convex function for all $i\in[1,\cdots,N]$.
Consider the functional
$\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})=\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}m_{i}(dx_{i})\Bigr)^{\frac{1}{\alpha_{i}}}$
with $\alpha_{i}=\frac{n+r_{i}}{np_{i}}$ on the set of even admissible tuples $\{V_{i}\}$. Let $\Phi_{1},\cdots,\Phi_{N}$ be a maximizing tuple of (9). Then every function $\Phi_{j}-\Phi_{j}(0)$ is $\beta_{j}$-homogeneous with $\beta_{j}=\Bigl(\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\Bigr)\bigl(1+\frac{r_{j}}{n}\bigr).$
In section 2 we discuss the relation of the Blaschke–Santaló inequality to the optimal transportation problem. We show in sections 4-5 that $(p_{1},\cdots,p_{N})$-homogeneous cost functions admit other nice properties. In particular, we prove that the corresponding functional and set versions of the Blaschke–Santaló inequality for these cost functions are equivalent. In addition, we show that the Blaschke–Santaló functional increases under the appropriate version of the Steiner symmetrization procedure.
Acknowledgements
The author was supported by RSF project 25-11-00007. The author thanks Matthieu Fradelizi for his interest and comments about the bibliography.
2. Optimal transportation approach
Let us remind the reader the basic facts about the mass transportation problem. Given $N$ probability measures $\mu_{i}$ on $\mathbb{R}^{n}$ and a cost function $c(x)=c(x_{1},\cdots,x_{n})$ we consider Kantorovich functional
$K^{\max}_{c}(\pi)=\int_{(\mathbb{R}^{n})^{N}}cd\pi,$
where $\pi$ is a measure on $(\mathbb{R}^{n})^{N}$ with marginal $\mu_{i}$: ${\rm Pr}_{i}\gamma=\mu_{i}$. Here we are interested in the maximization problem. If the value of $K^{\max}_{c}$ is maximized on some measure $\gamma$, we say that $\gamma$ is a solution to the Kantorovich maximization problem for the cost function $c$. For existence of $\gamma$ it is sufficient to assume that $c$ is upper-semicontinuous and admits some integrability properties (see [7], [37]).
The corresponding linear dual functional $D$ is defined on tuples of functions
$(f_{1}(x_{1}),\cdots,f_{N}(x_{N}))$
satisfying the assumption $\sum_{i=1}^{N}f_{i}(x_{i})\geq c$ and takes the form
$D(f_{1},\cdots,f_{N})=\sum_{i=1}^{N}\int_{\mathbb{R}^{n}}f_{i}d\mu_{i}.$
We always have
$K^{max}_{c}\leq D.$
Under broad assumptions on $c$ the dual functional $D$ admits minimum and this minimum coincides with the maximum of the Kantorovich functional (no duality gap). If the tuple
$(f_{1}(x_{1}),\cdots,f_{N}(x_{N}))$
is a minimum point of $D$, the complimentary slackness condition implies
$\sum_{i=1}f(x_{i})=c(x)$
for $\gamma$-almost all $x=(x_{1},\cdots,x_{N})$.
Without loss of generality we can always assume that our admissible tuple $\{f_{i}\}$ satisfies the following property:
$f_{i}(x_{i})=\sup_{x_{j},j\neq i}\Bigl(c(x)-\sum_{j\neq i}f(x_{j})\Bigr),$
i.e. every $f_{i}$ is the multiple $c$-Legendre transform of the other functions (see [34]).
In what follows we consider $N$-tuples of functions $V_{1}(x_{1}),\cdots,V_{N}(x_{N})$, $x_{i}\in\mathbb{R}^{n}$, satisfying assumption
$\sum_{i=1}^{N}V_{i}(x_{i})\geq c$
for a given cost function $c=c(x_{1},\cdots,x_{N})$ and the following Blaschke–Santaló-type functional
(16) $\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})=\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}m_{i}(dx_{i})\Bigr)^{\frac{1}{\alpha_{i}}}.$
Here $m_{i}$ are some probability reference measures on $\mathbb{R}^{n}$ and $\alpha_{i}$ are given positive numbers.
Remark 2.1.
In this section we don’t assume that $m_{i}$ and $c$ are homogeneous.
We associate with the tuple $\{V_{i}(x_{i})\}$ a tuple of probability measures
(17) $\nu_{i}=\frac{e^{-\alpha_{i}V_{i}}}{\int e^{-\alpha_{i}V_{i}}dm_{i}}\cdot m_{i},\ 1\leq i\leq N.$
The Blaschke–Santaló functional admits the following monotonicity property related to the multimarginal optimal transportation problem with marginals $\nu_{i}$. In a less general form it was first observed in [27].
Theorem 2.2.
(Monotonicity of the Blaschke–Santaló functional under optimal transportation). Assume that
  • the optimal transportation problem with marginals $\nu_{i}$ and the cost function $c$ admits a solution $\gamma\in\mathcal{P}((\mathbb{R}^{n})^{N})$
  • the corresponding dual transportation problem admits a solution $\Phi_{1},\cdots,\Phi_{N}$
  • no duality gap holds.
Then
$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}\Phi_{i}}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}.$
Proof.
By the assumption on $\{V_{i}\}$ and the complimentary slackness condition
$\sum_{i=1}^{N}\Phi_{i}(x_{i})=c(x)\leq\sum_{i=1}^{N}V_{i}(x_{i})$
for $\gamma$-almost all $(x_{1},\cdots,x_{N})$. Take
$\alpha=\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}$
and integrate inequality $1\leq e^{\alpha\sum_{i=1}^{N}{(V_{i}-\Phi_{i})}}$ over $\gamma$. By the Hölder inequality
$1\leq\int_{(\mathbb{R}^{n})^{N}}e^{\alpha\sum_{i=1}^{N}{(V_{i}-\Phi_{i})}}d\gamma\leq\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{\alpha_{i}(V_{i}-\Phi_{i})}d\gamma\Bigr)^{\frac{\alpha}{\alpha_{i}}}=\Bigl(\prod_{i=1}^{N}\frac{\int_{\mathbb{R}^{n}}e^{-\alpha_{i}\Phi_{i}}dm_{i}}{\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}}dm_{i}}\Bigr)^{\frac{\alpha}{\alpha_{i}}}.$
The result follows immediately from this inequality. ∎
Corollary 2.3.
Assume that the maximum of $\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})$ over all admissible tuples is reached on the tuple $\{\Phi_{i}\}$. Assume, in addition, that all the assumptions of the previous theorem are fulfilled for the cost $c$ and probability measures $\mu_{i}=\frac{e^{-\alpha_{i}\Phi_{i}}}{\int e^{-\alpha_{i}\Phi_{i}}dm_{i}}\cdot m_{i}$. Then $\{\Phi_{i}\}$ solves the dual transportation problem for $c$ and $\mu_{i}$.
The following result is a generalization of the sharp Talagrand inequality (5) (see [17], [22], [18], [15]). For $N>2$ it was known in less general situation (see [27]). Let $\{\Phi_{i}\}$ be maximizers of (16) and $\mu_{i}=\frac{e^{-\alpha_{i}\Phi_{i}}dm_{i}}{\int e^{-\alpha_{i}\Phi_{i}}dm_{i}}$. Introduce the positive cost function
$d(x)=\sum_{i=1}^{N}\Phi_{i}(x)-c(x).$
Given $N$ probability measures $\{\nu_{i}\},1\leq i\leq N,$ consider the multimarginal minimization transportation problem for $d$:
$K^{\min}_{d}(\nu_{1},\cdots,\nu_{N})=\inf_{\pi\in\Pi(\nu_{1},\cdots,\nu_{N})}\int d(x)d\pi,$
where $\Pi(\nu_{1},\cdots,\nu_{N})$ is the set of measures with projections $\nu_{1},\cdots,\nu_{N}$.
Theorem 2.4.
Assume that the maximum of $\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})$ over all admissible tuples is reached on the tuple $\{\Phi_{i}\}$. Assume that for a tuple of probability measures $\{\nu_{i}\}$ and the cost function $d(x)=\sum_{i=1}^{N}\Phi_{i}(x)-c(x)$ both the primal and the dual transportation problem admit solutions and no duality gap holds. Then
$K^{\min}_{d}(\nu_{1},\cdots,\nu_{N})\leq\sum_{i=1}^{N}\frac{1}{\alpha_{i}}{\rm Ent}_{\mu_{i}}(\nu_{i}).$
Proof.
Let $\{f_{i}(x_{i})\}$ be a solution to the dual transportation problem for $\{\nu_{i}\}$ and $d$. In particular,
$\sum_{i=1}^{N}f_{i}(x_{i})\leq\sum_{i=1}^{N}\Phi_{i}(x_{i})-c(x).$
Since $\sum_{i=1}^{N}(\Phi_{i}-f_{i})(x_{i})\geq c(x),$ one has
$\prod_{i=1}^{N}\Bigl(\int e^{\alpha_{i}(f_{i}-\Phi_{i})}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\prod_{i=1}^{N}\int e^{-\alpha_{i}\Phi_{i}}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}.$
by optimality of $\{\Phi_{i}\}$. This inequality can be rewritten as follows:
$\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\log\int e^{\alpha_{i}f_{i}}d\mu_{i}\leq 0.$
Let $\nu_{i}$ have densities with respect to $\mu_{i}$: $\nu_{i}=\rho_{i}\cdot\mu_{i}$ (otherwise the result is trivial). Applying inequality $xy\leq e^{x}-1+y\log y$ one gets the result:
$K_{d}(\nu_{1},\cdots,\nu_{N})$ $=\sum_{i=1}^{N}\int f_{i}d\nu_{i}\leq\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\int\Bigl(\alpha_{i}f_{i}-\log\int e^{\alpha_{i}f_{i}}d\mu_{i}\Bigr)\rho_{i}d\mu_{i}$
$\leq\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\Bigl(\rho_{i}\log\rho_{i}-1+e^{\alpha_{i}f_{i}-\log\int e^{\alpha_{i}f_{i}}d\mu_{i}}\Bigr)d\mu_{i}=\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\int\rho_{i}\log\rho_{i}d\mu_{i}$
$=\sum_{i=1}^{N}\frac{1}{\alpha_{i}}{\rm Ent}_{\mu_{i}}(\nu_{i}).$
Moreover, the following reverse Theorem is true.
Theorem 2.5.
Let $c$ be a cost function and $\{\Phi_{i}(x_{i})\}$ be a tuple of functions satisfying $\sum_{i=1}^{N}\Phi(x_{i})\geq c(x)$. Assume that the Kantorovich functional $K^{\min}_{d}$ with $d=\sum_{i=1}^{N}\Phi_{i}(x_{i})-c(x)$ satisfies inequality
$K^{\min}_{d}(\nu_{1},\cdots,\nu_{N})\leq\sum_{i=1}^{N}\frac{1}{\alpha_{i}}{\rm Ent}_{\mu_{i}}(\nu_{i})$
for $\mu_{i}=\frac{e^{-\alpha_{i}\Phi_{i}(x_{i})}dx_{i}}{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}$ and arbitrary tuple of probability measures $\{\nu_{i}\}$. Then $\{\Phi_{i}\}$ is the maximum point of $\mathcal{BS}_{\alpha,m}$.
Proof.
Let $\{V_{i}\}$ be a tuple of functions satisfying $\sum_{i}V_{i}\geq c.$ Define probability measures $\nu_{i}=\frac{e^{-\alpha_{i}V_{i}}dx_{i}}{\int e^{-\alpha_{i}V_{i}}dx_{i}}.$ One has ${\rm Ent}_{\mu_{i}}(\nu_{i})=\int(\alpha_{i}\Phi_{i}-\alpha_{i}V_{i})d\nu_{i}+\log\Bigl(\frac{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}{\int e^{-\alpha_{i}V_{i}}dx_{i}}\Bigr).$ Hence
$\sum_{i=1}^{N}\frac{1}{\alpha_{i}}{\rm Ent}_{\mu_{i}}(\nu_{i})=\sum_{i=1}^{N}(\Phi_{i}-V_{i})d\nu_{i}+\log\prod_{i=1}^{N}\Bigl(\frac{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}{\int e^{-\alpha_{i}V_{i}}dx_{i}}\Bigr)^{\frac{1}{\alpha_{i}}}.$
On the other hand $\sum_{i=1}^{N}(\Phi_{i}-V_{i})\leq\sum_{i=1}^{N}\Phi_{i}-c$. Hence by duality (the primal functional dominates the dual functional in the minimization problem)
$K^{\min}_{d}(\nu_{1},\cdots,\nu_{N})\geq\sum_{i=1}^{N}\int(\Phi_{i}-V_{i})d\nu_{i}.$
Finally, the assumption of the theorem implies
$\sum_{i=1}^{N}\int(\Phi_{i}-V_{i})d\nu_{i}\leq K^{\min}_{d}(\nu_{1},\cdots,\nu_{N})\leq\sum_{i=1}^{N}\frac{1}{\alpha_{i}}{\rm Ent}_{\mu_{i}}(\nu_{i})=\sum_{i=1}^{N}(\Phi_{i}-V_{i})d\nu_{i}+\log\prod_{i=1}^{N}\Bigl(\frac{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}{\int e^{-\alpha_{i}V_{i}}dx_{i}}\Bigr)^{\frac{1}{\alpha_{i}}}$
and one gets the desired inequality. ∎
3. Homogeneity of the maximizers
In this section we generalize Theorem B from [11] and find sufficient conditions for homogeneity of maximizers of $\mathcal{BS}_{\alpha,m}$. It will be always assumed that $m_{i}$ satisfy assumption (10):
$\rho_{i}(tx)=t^{r_{i}}\rho(x)$
for all $t\geq 0$ and $x_{i}\in\mathbb{R}^{n}$. The cost function $c$ is assumed to be at least $p$-homogeneous:
$c(tx)=t^{p}c(x)$
for all $t\geq 0$ and $x\in(\mathbb{R}^{n})^{N}$, but for the main result we ask for the stronger assumption of $(p_{1},\cdots,p_{N})$-homogeneity:
$c(x_{1},\cdots,t_{i}x_{i},\cdots,x_{N})=t^{p_{i}}_{i}c(x)$
for all $1\leq i\leq N,x\in(\mathbb{R}^{n})^{N},t_{i}\geq 0$.
Recall the standard notation
${\rm Var}_{\mu}f=\int f^{2}d\mu-\Bigl(\int fd\mu\Bigr)^{2}.$
Lemma 3.1.
Consider $\mathcal{BS}_{\alpha,m}$ functional, where $c$ is $p$-homogeneous and every $m_{i}$ satisfies (10). Let $\{\Phi_{i}\}$ be the tuple of maximizers for $\mathcal{BS}_{\alpha,m}$ and $\mu_{i}=\frac{e^{-\alpha_{i}\Phi_{i}}}{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}dx_{i}$. Then the following is true:
(18) $\int cd\gamma=\sum_{i=1}^{N}\int\Phi_{i}d\mu_{i}=\sum_{i=1}^{N}\frac{n+r_{i}}{p\alpha_{i}}$
(19) $\sum_{i=1}^{N}\alpha_{i}{\rm Var}_{\mu_{i}}(\Phi_{i})\leq\sum_{i=1}^{N}\frac{n+r_{i}}{p\alpha_{i}}$
(20) ${\rm Var}_{\gamma}(c)\leq\Bigl({\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\Bigr)\sum_{i=1}^{N}\alpha_{i}{\rm Var}_{\mu_{i}}(\Phi_{i}).$
In addition, one has equality in the last inequality if and only if
$\alpha_{i}\Bigl(\Phi_{i}(x_{i})-\int\Phi_{i}d\mu_{i}\Bigr)=\alpha_{j}\Bigl(\Phi_{j}(x_{j})-\int\Phi_{j}d\mu_{j}\Bigr)$
$\gamma$-almost everywhere.
Proof.
We observe that functions $\{\Phi_{i,t}(x_{i})=\frac{\Phi_{i}(tx_{i})}{t^{p}}\}$ satisfy the same constraint $\sum_{i=1}^{N}\Phi_{i,t}(x_{i})\geq c(x)$ for all $t>0$. Hence $f(t)=\mathcal{BS}_{\alpha,m}(\Phi_{1,t},\cdots,\Phi_{N,t})\leq\mathcal{BS}_{\alpha,m}(\Phi_{1},\cdots,\Phi_{N})$. Applying change of variables $tx_{i}=y_{i}$ and using homogeneity of $m_{i}$, one gets
$\mathcal{BS}_{\alpha,m}(\Phi_{1,t},\cdots,\Phi_{N,t})=\prod_{i=1}^{N}\Bigl(\int e^{-\frac{\alpha_{i}}{t^{p}}\Phi_{i}(tx_{i})}\rho_{i}(x_{i})dx_{i}\Bigr)^{\frac{1}{\alpha_{i}}}=\frac{1}{t^{\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}}}\prod_{i=1}^{N}\Bigl(\int e^{-\frac{\alpha_{i}}{t^{p}}\Phi_{i}(y_{i})}\rho_{i}(y_{i})dy_{i}\Bigr)^{\frac{1}{\alpha_{i}}}.$
Let $s=\frac{1}{t^{p}}$. The function
$g(s)=\log\mathcal{BS}_{\alpha,m}(\Phi_{1,s^{-\frac{1}{p}}},\cdots,\Phi_{N,s^{-\frac{1}{p}}})=\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\log\int e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy+\log s\sum_{i=1}^{N}\frac{n+r_{i}}{p\alpha_{i}}$
attains maximum at $s=1$. Hence
$g^{\prime}(s)=\frac{1}{s}\sum_{i=1}^{N}\frac{n+r_{i}}{p\alpha_{i}}-\sum_{i=1}^{N}\frac{\int\Phi_{i}e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy}{\int e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy},$
$g^{\prime\prime}(s)=-\frac{1}{s^{2}}\sum_{i=1}^{N}\frac{n+r_{i}}{p\alpha_{i}}+\sum_{i=1}^{N}\alpha_{i}\Bigl[\frac{\int\Phi^{2}_{i}e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy}{\int e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy}-\Bigl(\frac{\int\Phi_{i}e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy}{\int e^{-\alpha_{i}s\Phi_{i}(y)}\rho_{i}(y)dy}\Bigr)^{2}\Bigr].$
Finally, (18) follows from the first order condition $g^{\prime}(1)=0$ and (19) follows from the second order condition $g^{\prime\prime}(1)\leq 0$.
Let us show the third inequality. Define for brevity $\hat{\Phi}_{i}=\Phi_{i}-\int\Phi_{i}d\mu_{i}$. We apply the following identity
(21) $\sum_{i=1}^{N}\alpha_{i}(\hat{\Phi}_{i})^{2}=\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\Bigl[\bigl(\sum_{i=1}^{N}\hat{\Phi}_{i}\bigr)^{2}+\frac{1}{\alpha_{i}\alpha_{j}}\sum_{i\neq j}(\alpha_{i}\hat{\Phi}_{i}-\alpha_{j}\hat{\Phi}_{j})^{2}\Bigr].$
One has
$\sum_{i=1}^{N}\alpha_{i}\int(\hat{\Phi}_{i})^{2}d\mu_{i}=\int\sum_{i=1}^{N}\alpha_{i}(\hat{\Phi}_{i})^{2}d\gamma\geq\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\int\bigl(\sum_{i=1}^{N}\tilde{\Phi}_{i}\bigr)^{2}d\gamma=\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\Bigl[\int c^{2}d\gamma-\Bigl(\int cd\gamma\Bigr)^{2}\Bigr].$
Clearly, for the equality it is necessary and sufficient to have $\alpha_{i}\hat{\Phi}_{i}=\alpha_{j}\hat{\Phi}_{j}$ $\gamma$-a.e. ∎
Finally, the following theorem gives some sufficient conditions for homogeneity of the maximizing tuple.
Theorem 3.2.
Let $m_{i}$ satisfy assumption (10) and $c$ be $(p_{1},\cdots,p_{N})$-homogeneous. Consider functional (16) with
$\alpha_{i}=\frac{n+r_{i}}{np_{i}}.$
Let $\{\Phi_{i}\}$ be a maximizing tuple of (16). Assume, in addition, that
  • the cost function $c$ is locally Lipschitz
  • every function $\Phi_{j}$ is convex and $\Phi_{j}(0)<\infty$.
Then every function $\Phi_{j}-\Phi_{j}(0)$ is $\beta_{j}$-homogeneous with
$\beta_{j}=\Bigl(\sum_{i=1}^{N}\frac{1}{\alpha_{i}}\Bigr)\bigl(1+\frac{r_{j}}{n}\bigr).$
Remark 3.3.
Here we don’t recover the result from [31] on homogeneity of maximizers for $c_{bar}$, because the barycentric cost function is not $(p_{1},\cdots,p_{N})$-homogeneous.
Proof.
By Corollary 2.3 the tuple $\{\Phi_{i}\}$ is a solution to the dual multimarginal transportation problem for $c$ and $\mu_{i}=\frac{e^{-\alpha_{i}\Phi_{i}}}{\int e^{-\alpha_{i}\Phi_{i}}dx_{i}}.$ In particular, $\sum_{i=1}^{N}\Phi_{i}(x_{i})=c(x)$ for $\gamma$-almost all $x=(x_{1},\cdots,x_{N})$, where $\gamma$ is a solution to the primal transportation problem. In addition, since $\sum_{i=1}^{N}\Phi_{i}(x_{i})\geq c(x)$ for all $x$, it follows from the regularity assumptions on $c$ and $\Phi_{i}$ that
$\nabla\Phi_{i}(x_{i})=\nabla_{x_{i}}c(x)$
$\gamma$-a.e. Taking the scalar product of the latter identity, we get
(22) $\langle\nabla\Phi_{i}(x_{i}),x_{i}\rangle=\langle\nabla_{x_{i}}c(x),x_{i}\rangle=p_{i}c(x).$
According to Lemma 3.1 $\int cd\gamma=\frac{1}{\sum_{i=1}p_{i}}\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}=n.$ Applying (22) we get, in particular
$n=\int cd\gamma=\frac{1}{p_{i}}\int\langle x_{i},\nabla\Phi_{i}(x_{i})\rangle d\mu_{i}=\frac{1}{p_{i}\int_{\mathbb{R}^{n}}e^{-\alpha_{i}\Phi_{i}}\rho_{i}dx_{i}}\int_{\mathbb{R}^{n}}\langle x_{i},\nabla\Phi_{i}(x_{i})\rangle e^{-\alpha_{i}\Phi_{i}}\rho_{i}dx_{i}.$
Without loss of generality we assume that $\Phi_{i}$ is lower-semicontinuous and apply the following integration by parts formula
$\int_{\mathbb{R}^{n}}{\rm{div}}(v)e^{-\alpha_{i}\Phi_{i}}dx_{i}=\alpha_{i}\int_{\Omega}\langle v,\nabla\Phi_{i}\rangle e^{-\alpha_{i}\Phi_{i}}dx_{i}+\int_{\partial\Omega}\langle v,\nu\rangle e^{-\alpha_{i}\Phi_{i}}d\mathcal{H}^{n-1}.$
Here $\Omega=\{\Phi_{i}<\infty\}$, $\nu$ is the outer normal vector on $\partial{\Omega}$, $\mathcal{H}^{n-1}$ is the n1n-1-dimensional Haussdorff measure on $\partial\Omega$, and $v$ is a sufficiently regular vector field.
In particular, we obtain
$n=\frac{1}{p_{i}\alpha_{i}{\int_{\mathbb{R}^{n}}e^{-\alpha_{i}\Phi_{i}}\rho_{i}dx_{i}}}\Bigl[\int{\rm{div}}\bigl(\rho_{i}\cdot x_{i}\bigr)e^{-\alpha_{i}\Phi_{i}}dx_{i}-\int_{\partial\Omega}\langle x_{i},\nu\rangle e^{-\alpha_{i}\Phi_{i}}\rho_{i}d\mathcal{H}^{n-1}\Bigr].$
Using that ${\rm{div}}\bigl(\rho_{i}\cdot x_{i}\bigr)=(n+r_{i})\rho_{i}$, one gets $n=\frac{n+r_{i}}{p_{i}\alpha_{i}}-\frac{\int_{\partial\Omega}\langle x_{i},\nu\rangle e^{-\alpha_{i}\Phi_{i}}\rho_{i}d\mathcal{H}^{n-1}}{p_{i}\alpha_{i}{\int_{\mathbb{R}^{n}}e^{-\alpha_{i}\Phi_{i}}\rho_{i}dx_{i}}}.$ Since $\frac{n+r_{i}}{p_{i}\alpha_{i}}=n$, one finally gets
(23) $\int_{\partial\Omega}\langle x_{i},\nu\rangle e^{-\alpha_{i}\Phi_{i}}\rho_{i}d\mathcal{H}^{n-1}=0.$
Next we observe that $\langle x_{i},\nu\rangle\geq 0$, because $\Omega$ is convex as a sublevel set of the convex function $\Phi_{i}$ and contains the origin. This implies that either $\partial\Omega$ is empty or $\Phi_{i}|_{\partial\Omega=+\infty}$ $\mathcal{H}^{n-1}$-a.e. (equivalently, $\lim_{x\to y,x\in{\rm Int}(\Omega)}\Phi_{i}(x)=+\infty$ for $\mathcal{H}^{n-1}$-almost all $y\in\partial\Omega$).
Finally, one gets that the following integration by parts formula is valid:
(24) $\int_{\mathbb{R}^{n}}{\rm div}(v)e^{-\alpha_{i}\Phi_{i}}dx_{i}=\alpha_{i}\int_{\mathbb{R}^{n}}\langle v,\nabla\Phi_{i}\rangle e^{-\alpha_{i}\Phi_{i}}dx_{i}.$
Clearly, the formula works for vector fields which are sufficiently regular and integrable.
Next we apply another result of Lemma 3.1:
(25) $\sum_{i=1}^{N}\alpha_{i}{\rm Var}_{\mu_{i}}(\Phi_{i})\leq\frac{1}{\sum_{i=1}p_{i}}\sum_{i=1}^{N}\frac{n+r_{i}}{\alpha_{i}}=n.$
Using that $c=\sum_{i=1}^{N}\Phi_{i}=\langle x_{i},\nabla\Phi_{i}(x_{i})\rangle$ $\gamma$-a.e. and integration by parts formula (24), we compute $\int c^{2}d\gamma$:
$\int c^{2}d\gamma$ $=\sum_{i=1}^{n}\int c\Phi_{i}d\gamma=\sum_{i=1}^{n}\frac{1}{p_{i}}\int\langle x_{i}\nabla\Phi_{i}\rangle\Phi_{i}d\mu_{i}=\sum_{i=1}^{N}\frac{1}{p_{i}\alpha_{i}}\frac{\int{\rm div}(x_{i}\cdot\Phi_{i}\rho_{i})e^{-\alpha_{i}\Phi_{i}}dx_{i}}{\int e^{-\alpha_{i}\Phi_{i}}\rho_{i}dx_{i}}$
$=\sum_{i=1}^{N}\Bigl[\frac{n+r_{i}}{p_{i}\alpha_{i}}\int\Phi_{i}d\mu_{i}+\frac{1}{p_{i}\alpha_{i}}\int\langle x_{i},\nabla\Phi_{i}\rangle d\mu_{i}\Bigr]=n\sum_{i=1}^{N}\int\Phi_{i}d\mu_{i}+\bigl(\sum_{i=1}^{n}\frac{1}{\alpha_{i}}\bigr)\int cd\gamma$
$=n^{2}+\bigl(\sum_{i=1}^{n}\frac{1}{\alpha_{i}}\bigr)n.$
Thus we get $\int c^{2}d\gamma-\bigl(\int cd\gamma\bigr)^{2}=\bigl(\sum_{i=1}^{n}\frac{1}{\alpha_{i}}\bigr)n.$ Hence one gets from (21) that $\sum_{i=1}^{N}\alpha_{i}{\rm Var}_{\mu_{i}}(\Phi_{i})\geq n$, Finally, we get by the opposite inequality (25) that $\sum_{i=1}^{N}\alpha_{i}{\rm Var}_{\mu_{i}}(\Phi_{i})=n$. This can be possible only if $\alpha_{i}(\Phi_{i}(x_{i})-\int\Phi_{i}d\mu_{i})=\alpha_{j}(\Phi_{j}(x_{j})-\int\Phi_{j}d\mu_{j})$ $\gamma$-a.e. Since
$c(x)=\sum_{i=1}^{N}\Phi_{i}(x_{i})=\frac{1}{p_{j}}\langle x_{j},\nabla\Phi_{j}(x_{j})\rangle,$
we obtain that for $\mu_{j}$-almost all $x_{j}$ one has
$\frac{1}{p_{j}}\langle x_{j},\nabla\Phi_{j}(x_{j})\rangle=\Phi_{j}(x_{j})\Bigl(1+\sum_{i\neq j}\frac{\alpha_{j}}{\alpha_{i}}\Bigr)+C_{j}$
for some constant $C_{j}$. Taking $x_{j}=0$ we get that $C_{j}=-\Phi_{j}(0)\Bigl(1+\sum_{i\neq j}\frac{\alpha_{j}}{\alpha_{i}}\Bigr)$. Hence
$\Phi_{j}(x_{j})-\Phi_{j}(0)=\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\frac{1}{\alpha_{j}p_{j}}\langle x_{j},\nabla\Phi_{j}(x_{j})\rangle=\frac{1}{\sum_{i=1}^{N}\frac{1}{\alpha_{i}}}\frac{n}{n+r_{j}}\langle x_{j},\nabla\Phi_{j}(x_{j})\rangle.$
This completes the proof. ∎
Corollary 3.4.
Let $m_{i}$ satisfy assumption (10) and densities $\rho_{i}$ are positive almost everywhere. Consider functional (16) on the set of even functions and let $\{\Phi_{i}\}$ be a maximizing tuple. Assume that
  1. (1)
    $c$ is a continuous $(p_{1},\cdots,p_{N})$-homogeneous function for some numbers $p_{i}>0$
  2. (2)
    for every $1\leq i\leq N$ there exist numbers $s_{j}\in\{-1,1\}$, $j\neq i$ such that
    $c(x_{1},\cdots,x_{i-1},-x_{i},x_{i+1},\cdots,x_{N})=c(s_{1}x_{1},\cdots,s_{i-1}x_{i-1},x_{i},s_{i+1}x_{i+1},\cdots,s_{N}x_{N})$
  3. (3)
    $c$ is convex with respect to every variable, i.e. for all $i\in[1,\cdots,N]$ and $x_{j}\in\mathbb{R}^{n}$, $j\neq i$, the mapping
    $x\to c(x_{1},\cdots,x_{i-1},x,x_{i+1},x_{N})$
    is a convex function.
Then the conclusion of Theorem 3.2 holds.
Proof.
We have to prove that every $\Phi_{i}$ is convex (together with evenness, this clearly implies that $\Phi_{i}(0)=\min\Phi_{i}<+\infty$). Every $\Phi_{i}$ must satisfy
$\Phi_{i}(x_{i})=\tilde{\Phi}_{i}(x_{i})=\sup_{x_{j},j\neq i}\Bigl(c(x)-\sum_{j\neq i}\Phi_{j}(x_{j})\Bigr),$
otherwise replacing $\Phi_{i}$ by $\tilde{\Phi}_{i}$ we will increase the value of the functional. We observe that $\tilde{\Phi}_{i}$ is even, provided all $\Phi_{j}$, $j\neq i$ are even. Indeed, this follows immediately from the assumption (2) and evenness of $\Phi_{j}$, $j\neq i$. Clearly, $\Phi_{i}$ is convex as a supremum of convex functions. ∎
Remark 3.5.
Corollary 3.4 implies, in particular, Theorem B from [11] about homogeneity of the maximizers of the functional
$\mathcal{F}_{p,\Psi}(V)=\Bigl(\int_{\mathbb{R}^{n}}e^{-V}dx\Bigr)^{\frac{1}{p}}\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{p-1}V^{*}}\det D^{2}\Psi^{*}dy\Bigr)^{\frac{1}{q}}.$
Here $\Psi$ is a given convex $p$-homogeneous function. The functional is considered on the set of even convex functions. The proof from [11] relies on the application of the maximum principle for the corresponding Monge–Ampère equation.
Remark 3.6.
The convexity assumptions of Corollary 3.4 and Theorem 3.2 seems to be too strong. Eventually, the conclusion of Theorem 3.2 holds if $\Phi_{i}$ have the property: function $f_{i}(t)=\Phi_{i}(tx)$ is increasing on $0,+\infty)$ for all $x$. Equivalently, the sublevel sets of $\Phi_{i}$ are starshaped (star sets). We don’t prove any results of this type here.
In the rest of the section we discuss how to give an equivalent formulation of our problem on $\mathbb{S}^{n-1}$. Under assumptions of Corollary 3.4 (or Theorem 3.2) the maximization problem for $\mathcal{BS}_{\alpha,m}$ can have only homogeneous solution $\{\Phi_{i}\}$ (up to addition of the constants). Without loss of generality, we assume that our maximizing tuple satisfies
$\Phi_{i}(0)=0$
for all $1\leq i\leq N$. Let us represent every $\Phi_{i}$ in the form
$\Phi_{i}(t\theta_{i})=\frac{p_{i}}{\beta_{i}}t^{\beta_{i}}\varphi^{\beta_{i}}_{i}(\theta_{i}$
for some positive function $\varphi_{i}$ on $\mathbb{S}^{n-1}$.
Lemma 3.7.
Admissibility of the tuple $\{\Phi_{i}\}$ is equivalent to the property
$c(\theta_{1},\cdots,\theta_{N})\leq\prod_{i=1}^{N}\varphi_{i}^{p_{i}}(\theta_{i}).$
Proof.
Assume that $c(\theta_{1},\cdots,\theta_{N})\leq\prod_{i=1}^{N}\varphi_{i}^{p_{i}}(\theta_{i})$, equivalently $c\Bigl(\frac{\theta_{1}}{\varphi_{1}(\theta_{1})},\cdots,\frac{\theta_{N}}{\varphi_{N}(\theta_{N})}\Bigr)\leq 1.$ We get by the Hölder inequality
$c(x)=\prod_{i=1}^{N}r_{i}^{p_{i}}\cdot c(\theta_{1},\cdots,\theta_{N})\leq\prod_{i=1}^{N}\varphi_{i}^{p_{i}}(\theta_{i})r_{i}^{p_{i}}\leq\sum_{i=1}^{N}\frac{p_{i}}{\beta_{i}}\varphi_{i}^{\beta_{i}}(\theta_{i})r_{i}^{\beta_{i}}=\sum_{i=1}^{N}\Phi_{i}(x_{i}),$
thus the tuple $\{\Phi_{i}\}$ is admissible.
If $\{\Phi_{i}\}$ is admissible, then $c\Bigl(\frac{\theta_{1}}{\varphi_{1}(\theta_{1})},\cdots,\frac{\theta_{N}}{\varphi_{N}(\theta_{N})}\Bigr)\leq\sum_{i=1}^{N}\frac{p_{i}}{\beta_{i}}\varphi_{i}^{\beta_{i}}(\theta_{i})\Bigl(\frac{1}{\varphi_{i}(\theta_{i})}\Bigr)^{\beta_{i}}=\sum_{i=1}^{N}\frac{p_{i}}{\beta_{i}}=1.$ The proof is complete. ∎
Functional $\mathcal{BS}_{\alpha,m}$ restricted to homogeneous functions $\{V_{i}\}$ can be rewritten in terms of functions $\{\varphi_{i}\}$.
$\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})$ $=\prod_{i=1}^{N}\bigl(\int_{\mathbb{S}^{n-1}}e^{-{\alpha_{i}}V_{i}(x_{i})}\rho_{i}(x_{i})dx_{i}\bigr)^{\frac{1}{\alpha_{i}}}$
$=\prod_{i=1}^{N}\Bigl(\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}e^{-\frac{\alpha_{i}p_{i}}{\beta_{i}}\varphi_{i}^{\beta_{i}}(\theta_{i})t^{\beta_{i}}_{i}}\rho_{i}(\theta_{i})t_{i}^{n-1+r_{i}}dt_{i}d\theta_{i}\Bigr)^{\frac{1}{\alpha_{i}}}$
Making change of variables $s_{i}=t_{i}\varphi_{i}(\theta_{i})$ we conclude that
$\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})=C\prod_{i=1}^{N}\Bigl(\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}\Bigr)^{\frac{1}{\alpha_{i}}},$
where $C$ is a constant, depending on parameters $\alpha_{i},p_{i},r_{i},n$.
As a conclusion, we obtain the following statement.
Proposition 3.8.
Under the above assumptions the problem of maximization of $\mathcal{BS}_{\alpha,m}$ is equivalent to maximization of the spherical Blaschke–Santaló type functional
(26) $\prod_{i=1}^{N}\Bigl(\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}\Bigr)^{\frac{1}{\alpha_{i}}}$
under constraints
(27) $c(\theta_{1},\cdots,\theta_{N})\leq\prod_{i=1}^{N}\varphi_{i}^{p_{i}}(\theta_{i}).$
Remark that the latter assumption can be rewritten in additive (transportation-type) form:
$\log c(\theta_{1},\cdots,\theta_{N})\leq\sum_{i=1}^{N}p_{i}\log\varphi_{i}(\theta_{i}),$
which makes this spherical problem more similar to the initial problem. See [25], [18] and the reference therein for the study of the logarithmic cost function
$\log\langle x,y\rangle$
in the case $N=2$.
Finally, the following Theorem generalizes Theorem 2.2 and Corollary 2.3.
Theorem 3.9.
Let $c\geq 0$ be sufficiently regular cost function on $(\mathbb{S}^{n-1})^{N}$, $\varphi_{i}$ be nonnegative functions satisfying (27) and, in addition, $\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}<\infty$ for all $1\leq i\leq N$. Consider mass transportation problem with cost function $\log c$ and probability measures
$\mu_{i}=\frac{1}{\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}.$
Let $\{u_{i}\},1\leq i\leq N$ be the solution to the corresponding dual problem. Define $\psi_{i}=e^{\frac{1}{p_{i}}u_{i}}$. If no duality gap for the cost $\log c$ holds, the following inequality is true
$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\prod_{i=1}^{N}\Bigl(\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\psi^{n+r_{i}}_{i}(\theta_{i})}\Bigr)^{\frac{1}{\alpha_{i}}}.$
In particular, if $\{\varphi_{i}\}$ is the maximum point for (26), then $\varphi_{i}=c_{i}e^{\frac{1}{p_{i}}u_{i}}$, where constants $c_{i}>0$ satisfy $\prod_{i=1}^{N}c_{i}=1$.
Proof.
Let $\gamma$ be the solution to the primal transportation problem. One has
$\log c=\sum_{i=1}^{N}u_{i}=\sum_{i=1}^{N}p_{i}\log\psi_{i}$
$\gamma$-a.e. Hence, for $\gamma$-almost all $(\theta_{1},\cdots,\theta_{N})$ one has $\prod_{i=1}^{N}\psi_{i}^{p_{i}}(\theta_{i})\leq\prod_{i=1}^{N}\varphi_{i}^{p_{i}}(\theta_{i}).$ Hence by the Hölder inequality
$1\leq\int\prod_{i=1}^{N}\frac{\varphi_{i}^{p_{i}}(\theta_{i})}{\psi_{i}^{p_{i}}(\theta_{i})}d\gamma\leq\prod_{i=1}^{N}\Bigl(\int\frac{\varphi_{i}^{n+r_{i}}(\theta_{i})}{\psi_{i}^{n+r_{i}}(\theta_{i})}d\gamma\Bigr)^{\frac{p_{i}}{n+r_{i}}}=\prod_{i=1}^{N}\Bigl(\frac{\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\psi^{n+r_{i}}_{i}(\theta_{i})}}{\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}}\Bigr)^{\frac{1}{n\alpha_{i}}}.$
This completes the proof. ∎
Remark 3.10.
The typical example of $c$ is $\langle x,y\rangle$ and the cost $\log c$ is defined to be $\infty$ outside of the area $\langle x,y\rangle>0$. Thus the questions of existence/duality are not obvious. However, everything works in the symmetric setting.
3.1. The classical Blaschke–Santaló inequality revisited
Let us quickly discuss transportation proofs of the classical Blaschke–Santaló inequality. By the results of this section the maximizer $\Phi$ of the functional
$\int_{\mathbb{R}^{n}}e^{-V}dx\int_{\mathbb{R}^{n}}e^{-V^{*}}dy$
(if exists) has the form
$\Phi=\frac{1}{2}\varphi^{2}(\theta)|x|^{2}.$
This was proved in [11], but Theorem 3.2 provides more soft arguments. The existence of a maximizer can be proved by compactness arguments (see [11]). By Corollary 2.3 $\nabla\Phi$ pushes forward $\mu=\frac{e^{-\Phi}}{\int e^{-\Phi}dx}$ onto $\nu=\frac{e^{-\Phi^{*}}}{\int e^{-\Phi^{*}}dx}$ and satisfies
(28) $\frac{e^{-\Phi}}{\int e^{-\Phi}dx}=\frac{e^{-\Phi^{*}}(\nabla\Phi)}{\int e^{-\Phi^{*}}dx}\det D^{2}\Phi.$
Since $\Phi^{*}(\nabla\Phi)=\Phi$, we also conclude that $\det{D^{2}\Phi}$ is constant. Using Theorem 3.9 one can associate with $\varphi$ Monge–Ampère equation on $\mathbb{S}^{n-1}$ (see [8], [11]). Finally, we can conclude that $\Phi$ is a quadratic function in several ways: 1) using the Pogorelov theorem describing global solutions to the equation $\det{D^{2}\Phi}=1$, 2) main result of [10] about solutions to (28), 3) the result on uniqueness (up to linear transformations) for the corresponding Monge–Ampère equation on the sphere.
4. Inequalities for sets and for functions
In this section we continue our study of homogeneous solutions and find a direct link between the functional and the set inequality. Let us specify assumptions on $c$ in this section.
Assumption on $c$:
  • $c$ is convex for every $x_{i}$, i.e.
    $x\to c(x_{1},\cdots,x_{i-1},x,x_{i+1},\cdots,x_{N})$
    is convex
  • $c$ is $p$-homogeneous for some $p>0$:
    $c(tx)=t^{p}c(x),\ t\geq 0.$
  • for every $1\leq i\leq N$ there exist numbers $s_{j}\in\{-1,1\}$, $j\neq i$ such that
    $c(x_{1},\cdots,x_{i-1},-x_{i},x_{i+1},\cdots,x_{N})=c(s_{1}x_{1},\cdots,s_{i-1}x_{i-1},x_{i},s_{i+1}x_{i+1},\cdots,s_{N}x_{N})$
We have already seen in the previous section that under these assumptions the multiple Legendre transform preserves convex even functions, more precisely function
$f_{i}(x_{i})=\sup_{x_{j},j\neq i}\bigl(c(x)-\sum_{j\neq i}f_{j}(x_{j})\bigr)$
is convex. Moreover, $f_{i}$ is even provided all $f_{j}$ are even, $j\neq i$.
We consider Blaschke–Santaló on the tuples of even functions. Since Lagrange transform preserves evenness and convexity, without loss of generality one can consider only convex even tuples $\{V_{i}\}$.
Remark 4.1.
Note that $\sum_{i=1}^{N}V_{i}(0)\geq c(0)=0$. Since 0 is the minimum point for every $V_{i}$, without loss of generality we can assume that $V_{i}\geq 0$ and $V_{i}(0)=0$.
Our aim is to relate the $\mathcal{BS}_{\alpha,m}$ to the following Blaschke–Santaló functional for sets:
$\mathcal{BS}^{set}_{m,\alpha}(K_{1},\cdots,K_{N})=\prod_{i=1}^{N}m_{i}(K_{i})^{\frac{1}{\alpha_{i}}}.$
Similarly, we consider only symmetric and convex sets. The $c$-polar transform
$K_{i}=\{x_{i}:c(x)\leq 1\ \ {\rm if}\ \ x_{j}\in K_{j},j\neq i\}$
preserves symmetric convex sets as well.
We are interested in the following set volume product maximization problem: find the maximum of $\mathcal{BS}^{set}_{m,\alpha}$ on tuples of even convex sets $\{K_{i}\}$ satisfying
(29) $c(x)\leq 1$
for all $x\in\prod_{i=1}^{N}K_{i}$.
In the following Proposition and Corollary we prove a generalization of the well-known result about equivalence of the set and the functional versions of the classical Blaschke–Santaló inequality (see [4], [1]).
Proposition 4.2.
Let $\{K_{i}\},1\leq i\leq N,$ be a solution to the set maximization problem. Then for every admissible tuple $\{V_{i}\}$ one has
$\mathcal{BS}_{\alpha,m}(V_{1},\cdots,V_{N})\leq\mathcal{BS}_{\alpha,m}(\tau_{1}\|x_{1}\|^{\beta_{1}}_{K_{1}},\cdots,\tau_{N}\|x_{N}\|^{\beta_{N}}_{K_{N}}),$
where $\|\cdot\|_{K_{i}}$ is the Minkowski functional of $K_{i}$ and $\tau_{i}=\frac{\frac{1}{\alpha_{i}}}{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}$.
Proof.
By the “layer-cake” representation, Prekopa–Leindler and Hölder inequalities one has
$\Bigl(\prod_{i=1}^{N}\int e^{-\alpha_{i}V_{i}(x_{i})}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}$ $=\Bigl(\prod_{i=1}^{N}\int_{0}^{\infty}m_{i}(\{x_{i}:\alpha_{i}V_{i}(x_{i})\leq t_{i}\})e^{-t_{i}}dt_{i}\Bigr)^{\tau_{i}\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}$
$\leq\Bigl[\int_{0}^{\infty}\sup_{t=\sum_{i}\tau_{i}t_{i}}\Bigl(\prod_{i=1}^{N}m_{i}(\{x_{i}:\alpha_{i}V_{i}(x_{i})\leq t_{i}\})\Bigr)^{\tau_{i}}e^{-t}dt\Bigr]^{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}$
$\leq\Bigl[\int_{0}^{\infty}\sup_{t=\sum_{i}\tau_{i}t_{i}}\Bigl(\prod_{i=1}^{N}m^{\frac{1}{\alpha_{i}}}_{i}(\tilde{A}_{i})\Bigr)^{\frac{1}{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}}e^{-t}dt\Bigr]^{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}},$
where $\tilde{A}_{i}=\Bigl\{x_{i}:V_{i}(x_{i})\leq\frac{t_{i}}{\alpha_{i}}\Bigr\}.$ One has $c(x)\leq\sum_{i}V_{i}(x_{i})\leq\sum_{i}\frac{t_{i}}{\alpha_{i}}$ for all $\{x_{i}\}$ such that $x_{i}\in\tilde{A}_{i}$. Set
$A_{i}=\frac{1}{(\sum_{j=1}^{N}\frac{t_{j}}{\alpha_{j}})^{\frac{1}{p}}}\tilde{A}_{i}.$
Using $p$-homogeneity of $c$ one gets $c(x)\leq 1$ provided $x_{i}\in{A}_{i}$. Applying inequality $\prod_{i=1}^{N}m_{i}(A_{i})^{\frac{1}{\alpha_{i}}}\leq\prod_{i=1}^{N}m_{i}(K_{i})^{\frac{1}{\alpha_{i}}}$ we obtain the following:
$\prod_{i=1}^{N}m_{i}(\tilde{A}_{i})^{\frac{1}{\alpha_{i}}}$ $=\Bigl(\sum_{j=1}^{N}\frac{t_{j}}{\alpha_{j}}\Bigr)^{\frac{1}{p}\sum_{i}\frac{n+r_{i}}{\alpha_{i}}}\prod_{i=1}^{N}m_{i}^{\frac{1}{\alpha_{i}}}({A}_{i})\leq\Bigl(\sum_{j=1}^{N}\frac{t_{j}}{\alpha_{j}}\Bigr)^{\frac{1}{p}\sum_{i}\frac{n+r_{i}}{\alpha_{i}}}\prod_{i=1}^{N}m_{i}^{\frac{1}{\alpha_{i}}}({K}_{i})$
$=\Bigl(\sum_{j=1}^{N}\frac{t_{j}}{\alpha_{j}}\Bigr)^{\sum_{i}\frac{n+r_{i}}{\alpha_{i}\beta_{i}}}\prod_{i=1}^{N}m_{i}^{\frac{1}{\alpha_{i}}}({K}_{i})=\prod_{i=1}^{N}m_{i}^{\frac{1}{\alpha_{i}}}\Bigl(\bigl(\sum_{j=1}^{N}\frac{t_{j}}{\alpha_{j}}\bigr)^{\frac{1}{\beta_{i}}}K_{i}\Bigr).$
Hence
$\int_{0}^{\infty}$ $\sup_{t=\sum_{i}\tau_{i}t_{i}}\Bigl(\prod_{i=1}^{N}m^{\frac{1}{\alpha_{i}}}_{i}(\tilde{A}_{i})\Bigr)^{\frac{1}{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}}e^{-t}dt\leq\int_{0}^{\infty}\Bigl(\prod_{i=1}^{N}m^{\frac{1}{\alpha_{i}}}_{i}\Bigl(\Bigl[t\sum_{j=1}^{N}\frac{1}{\alpha_{j}}\Bigr]^{\frac{1}{\beta_{i}}}K_{i}\Bigr)\Bigr)^{\frac{1}{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}}e^{-t}dt$
$\leq\prod_{i=1}^{N}\Bigl(\int_{0}^{\infty}m_{i}\Bigl(\Bigl[t\sum_{j=1}^{N}\frac{1}{\alpha_{j}}\Bigr]^{\frac{1}{\beta_{i}}}K_{i}\Bigr)e^{-t}dt\Bigr)^{\tau_{i}}=\prod_{i=1}^{N}\Bigl(\int_{0}^{\infty}m_{i}\Bigl(\|x_{i}\|^{\beta_{i}}_{K_{i}}\leq t\sum_{j=1}^{N}\frac{1}{\alpha_{j}}\Bigr)e^{-t}dt\Bigr)^{\tau_{i}}$
$=\prod_{i=1}^{N}\Bigl(\int e^{-\frac{\|x_{i}\|^{\beta_{i}}}{\sum_{j=1}^{N}\frac{1}{\alpha_{j}}}}dm_{i}\Bigr)^{\tau_{i}}.$
Finally, resuming all the inequalities above, one gets
$\Bigl(\prod_{i=1}^{N}\int e^{-\alpha_{i}V_{i}(x_{i})}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\prod_{i=1}^{N}\Bigl(\int e^{-{\alpha_{i}\tau_{i}\|x_{i}\|^{\beta_{i}}}}dm_{i}\Bigr)^{\frac{1}{\alpha_{i}}}.$
The proof is complete. ∎
Corollary 4.3.
If $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$ is admissible tuple, i.e. $c(x)\leq\sum_{i=1}^{N}\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}},$ then it is a maximizing tuple.
We observe that in general the tuple of functions functions $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$ can be not admissible, hence Proposition 4.2 does not imply in general that $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$ is a maximizing tuple.
We give below some sufficient conditions for the tuple to be admissible.
Theorem 4.4.
Let $c$ satisfy all the assumptions from the beginning of the section. Assume, in addition, that $c$ is $(p_{1},\cdots,p_{N})$-homogeneous,
$\alpha_{i}=\frac{1}{p_{i}}\bigl(1+\frac{r_{i}}{n}\bigr),\ \ \ \ \ \\ \beta_{i}=\bigl(1+\frac{r_{i}}{n}\bigr)\sum_{j=1}^{N}\frac{1}{\alpha_{j}}=\frac{p_{i}}{\tau_{i}}.$
Assume that $K_{1},\cdots,K_{N}$ is the maximizing tuple for $\mathcal{B}S^{set}_{m,\alpha}$. Then $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$ is admissible. Hence $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$ is the maximum point of $\mathcal{BS}_{\alpha,m}$ on the set of even convex functions.
Proof.
According to Lemma 4.3 it is sufficient to prove admissibility of $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$. Take arbitrary point $x$ such that $x_{i}\neq 0,1\leq i\leq N$. Then $\frac{x_{i}}{\|x_{i}\|_{K_{i}}}\in K_{i}$. Hence
$c\bigl(\frac{x_{1}}{\|x_{1}\|_{K_{1}}},\cdots,\frac{x_{N}}{\|x_{N}\|_{K_{N}}}\bigr)\leq 1.$
Using $p_{i}$-homogeneity and the Hölder inequality, one gets
$c(x)\leq\|x_{1}\|^{p_{1}}_{K_{1}}\cdots\|x_{N}\|^{p_{N}}_{K_{N}}\leq\sum_{i=1}^{N}\frac{p_{i}}{\beta_{i}}\|x_{i}\|^{\beta_{i}}_{K_{i}}=\sum_{i=1}^{N}\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}.$
Theorem 4.4 and Corollary 3.4 imply the following result.
Corollary 4.5.
Let all the assumptions of Corollary 3.4 be fulfilled. Then every maximizing tuple $\{\Phi_{i}\}$ of the functional $\mathcal{BS}_{\alpha,m}$ with $\Phi_{i}(0)=0$ has the form $\{\tau_{i}\|x_{i}\|^{\beta_{i}}_{K_{i}}\}$, where $\{K_{i}\}$ is the maximizing tuple of the functional $\mathcal{BS}^{set}_{\alpha,m}$.
Remark 4.6.
Let all the assumptions of Corollary 3.4 be fulfilled. Then the equivalent functional problem on $\mathbb{S}^{n-1}$ is described in Proposition 3.8.
5. Symmetrization
In this section we consider a generalized symmetrization procedure for $N>2$ functions possessing property to increase the Blaschke–Santaló functional.
Let $c(x_{1},\cdots,x_{N})$ be a continuous function of $N$ variables $x_{i}\in\mathbb{R}^{n}$. As everywhere in this paper we say that a tuple of convex symmetric sets $\{A_{i}\},1\leq i\leq N,A_{i}\subset\mathbb{R}^{n}$ is admissible if
$c(x_{1},\cdots,x_{N})\leq 1$
provided $x_{i}\in A_{i}$ for all $1\leq i\leq N$.
We will assume in this section that the polar $c$-transform preserves symmetric and convex sets, i.e.
$B_{j}=\{x_{j}:c(x_{1},\cdots,x_{N})\leq 1,\forall x_{i}\in B_{i},i\neq j\}$
is symmetric (convex) if $B_{i}$ are symmetric (convex) for all $j\neq i$. Recall that this is true if $c$ satisfies assumption (2) of Corollary 3.4.
Let us describe the symmetrization with respect to a coordinate unit vector $e_{j}$, $1\leq j\leq n$ for a given couple of indices $i_{1}\neq i_{2}$, $1\leq i_{1},i_{2}\leq N$. We will call this procedure $(j,i_{1},i_{2})$-symmetrization. The procedure will replace sets $A_{i_{1}},A_{i_{2}}$ by some sets $B_{i_{1}},B_{i_{2}}$ and does not change all the other sets $A_{i}$.
Given a coordinate unit vector $e_{j}$ we represent arbitrary $x_{i}\in\mathbb{R}^{n}$ in the form
$x_{i}=(y_{i},t_{i})=y_{i}+t_{i}e_{j},\ \ \ y_{i}\bot e_{j}.$
Definition 5.1.
We say that the cost function $c$ satisfies $(j,i_{1},i_{2})$-assumption if the following holds:
  • The function
    $(y_{i_{1}},t_{i_{2}})\to c((y_{1},t_{1}),\cdots,(y_{N},t_{N}))$
    is convex for all fixed $y_{i_{k}}$, $k\neq 1$ and $t_{i_{j}}$, $j\neq 2$
  • $c((y_{1},t_{1}),$ $\cdots,(y_{i_{1}},-t_{i_{1}}),\cdots,(y_{i_{2}},t_{i_{2}}),\cdots(y_{N},t_{N}))$
    $=c((y_{1},t_{1}),\cdots,(y_{i_{1}},t_{i_{1}}),\cdots,(y_{i_{2}},-t_{i_{2}}),\cdots(y_{N},t_{N})).$
In what follows we write
$x_{i}=(x_{i,1},\cdots,x_{i,n})=(x_{i,j}),1\leq j\leq n,$
for every $1\leq i\leq N$.
Example 5.2.
Function
$c_{n}(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}=\sum_{j=1}^{n}\prod_{i=1}^{N}x_{i,j}$
admits the following property: $c_{n}((y_{1},t_{1}),\cdots,(y_{N},t_{N}))=c_{n-1}(y_{1},\cdots,y_{N})+\prod_{i=1}^{N}t_{i}.$ Clearly, $c$ satisfies $(j,i_{1},i_{2})$-assumption for all $j,i_{1},i_{2}$.
In addition, function $\sum_{j=1}^{n}\prod_{i=1}^{N}|x_{i,j}|^{p_{ij}}$ satisfies $(j,i_{1},i_{2})$-assumption for all $j,i_{1},i_{2}$ if $p_{i,j}\geq 1$ for all $i,j$.
Recall that given a convex set $A$, the set $\tilde{A}$ obtained by the Steiner symmetrization with respect to $e_{j}$ can be represented as follows:
$\tilde{A}=\Bigl\{\big(y,\frac{t-s}{2}\big),(y,t),(y,s)\in A\Bigr\}.$
Lemma 5.3.
Let $\{A_{i}\}$ be admissible tuple of symmetric convex bodies, where the cost function $c$ satisfies $(n,1,2)$-assumption and the polar $c$-transform preserves symmetric convex sets.
Let $m_{1}=\rho_{1}dx_{1},m_{2}=\rho_{2}dx_{2}$ be locally finite measures admitting the following properties:
  1. (1)
    $t\to\rho_{1}(y+te_{n})$ is an even function, decreasing on $0,+\infty)$ for all $y\bot e_{n}$
  2. (2)
    $y\to\rho_{2}(y+te_{n})$ is a symmetric log-concave function for all $t\in\mathbb{R}$.
Then there exists an admissible tuple of the form $\{\tilde{A}_{1},B,A_{3},\cdots,A_{N}\}$ such that
$m_{1}(\tilde{A}_{1})\geq m_{1}(A_{1}),\ \ \ m_{2}(B)\geq m_{2}(A_{2}).$
Proof.
We observe first that the Steiner symmetrization $\tilde{A}_{1}$ of $A_{1}$ satisfies $m_{1}(\tilde{A}_{1})\geq m_{1}(A_{1})$, this is a known property (see, for instance, Proposition 4.7 in [11] for the functional version of the statement and the references therein).
Without loss of generality one can take for $A_{2}$ the polar $c$-transform of $A_{1},A_{3},\cdots,A_{N}$:
$A_{2}=\{x_{2}:c(x)\leq 1,\mbox{for all }x_{i}\in A_{i},i\neq 2\}.$
Similarly we define $B$ as the $c$-transform of $\tilde{A}_{1}$ and $A_{i}$ with $i>2$:
$B=\{x_{2}:c(x)\leq 1,\mbox{for all }x_{1}\in\tilde{A}_{1},x_{i}\in A_{i},i>2\}.$
Since $c$-polar transform preserves symmetric convex bodies, both bodies are symmetric and convex.
For arbitrary $A\subset\mathbb{R}^{n}$ we denote by $A(r)$ the section of $A$ at the level $r$:
$A(r)=\{y:(y,r)=y+re_{n}\in A\}.$
We want to prove that
(30) $\frac{A_{2}(r)+A_{2}(-r)}{2}\subset B(r).$
From (30) we extract using Prekopa–Leindler inequality and assumptions of symmetry of $A_{2}$ and the log-concavity of $\rho_{2}(y+te)$:
$m_{2}(B)$ $=\int_{\mathbb{R}}\Bigl(\int_{B(r)}\rho_{2}(y+te_{n})dy\Bigr)dt\geq\int_{\mathbb{R}}\Bigl(\int_{\frac{A_{2}(r)+A_{2}(-r)}{2}}\rho_{2}(y+te_{n})dy\Bigr)dt$
$\geq\int_{\mathbb{R}}\Bigl(\int_{A_{2}(r)}\rho_{2}(y+te_{n})dy\Bigr)^{\frac{1}{2}}\Bigl(\int_{A_{2}(-r)}\rho_{2}(y+te_{n})dy\Bigr)^{\frac{1}{2}}dt=\int_{\mathbb{R}}\int_{A_{2}(r)}\rho_{2}(y+te_{n})dydt$
$=m_{2}(A_{2}).$
It remains to prove (30). One has
$B(r)=$ $\{v:c\bigl((y_{1},\frac{t-s}{2}),(v,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\leq 1,$
$\mbox{for all }(y_{1},t)\in A_{1},(y_{1},s)\in A_{1},(y_{i},t_{i})\in A_{i},i>2\}.$
Take arbitrary $a\in A_{2}(r),b\in A_{2}(-r)$. One has
$c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\leq 1,\ \ \ c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\leq 1,$
for all $(y_{1},t),(y_{1},s)\in A_{1},(y_{i},t_{i})\in A_{i},i>2$. Hence
$1$ $\geq\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
$=\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},-s),(b,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
$\geq c\bigl((y_{1},\frac{t-s}{2}),(\frac{a+b}{2},r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr).$
Thus $\frac{a+b}{2}\in B(r)$ and the proof is completed. ∎
Recall that a set $A$ is called unconditional with respect to the basis $e_{1},\cdots,e_{n}$ if
$x=\sum_{i=1}^{n}x_{i}e_{i}\in A\Longleftrightarrow\sum_{i=1}^{n}s_{i}x_{i}e_{i}\in A$
for any tuple $\{s_{i}\}$ such that $s_{i}\in\{-1,1\}$.
Corollary 5.4.
Let $\{m_{i}\}$ be log-concave unconditional measures, $1\leq i\leq N$ and $c$ satisfies $(j,i_{1},i_{2})$-assumption for all $j,i_{1},i_{2}$.
Assume, in addition, that the polar $c$-transform preserves 1) symmetric sets, 2) unconditional sets, 3) convex sets. Then for every admissible tuple $\{A_{i}\}$ of symmetric convex bodies there exists an admissible tuple $\{B_{i}\}$ of convex unconditional bodies with the property
$m_{i}(A_{i})\leq m_{i}(B_{i}).$
Proof.
We consequently apply symmetrization with respect to $e_{1},\cdots,e_{n}$ to the set $A_{1}$, keeping $A_{3},\cdots,A_{N}$ fixed and defining the second body in the tuple by the $c$-polar transform. By Lemma 5.3 one finally obtains another tuple of even convex sets $\{A^{\prime}_{i}\}$ such that $m_{i}(A^{\prime}_{i})\geq m_{i}(A_{i})$ and, in addition, $A^{\prime}_{1}$ is unconditional. Repeating this procedure N1N-1 times, we get unconditional sets $B_{1},\cdots,B_{N-1}$ and $B_{N}$ will be $c$-polar transform of $B_{1},\cdots,B_{N-1}$. Since this operation preserves unconditional sets, we get the claim. ∎
6. Main example
Proposition 6.1.
Let $f_{i}\colon\mathbb{R}_{+}^{n}\to\mathbb{R}_{+}$, $1\leq i\leq N$ be Borel functions, $\{\alpha_{i}\}$ be $N$ positive numbers and
$c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x^{\frac{1}{\alpha_{1}}}_{1,j}\cdots x^{\frac{1}{\alpha_{N}}}_{N,j}.$
Assume that functions $\{f_{i}(x_{i})\}$ satisfy inequality
$\prod_{i=1}^{N}f^{\frac{1}{\alpha_{i}}}_{i}(x_{i})\leq\rho(c(x))$
for some Borel function $\rho$. Then
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(x_{i})dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\bigl(\sum_{j=1}^{n}t^{A}_{j}\bigr)dt\Bigr)^{A},$
where $A=\sum_{j=1}^{N}\frac{1}{\alpha_{j}}$.
In particular, if functions $V_{i}(x_{i})$ satisfy $\sum_{i=1}^{N}V_{i}(x_{i})\geq\sum_{j=1}^{n}x^{\frac{1}{\alpha_{1}}}_{1,j}\cdots x^{\frac{1}{\alpha_{N}}}_{N,j}$, then
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}e^{-\frac{1}{A}\sum_{j=1}^{n}t^{A}_{j}}dt\Bigr)^{A}.$
Proof.
Following [19] we apply the exponential change of variables : $x_{i,j}=e^{t_{i,j}}$ (or simply $x_{i}=e^{t_{i}}$) and the Prekopa–Leindler inequality
$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})dx_{i,1}\cdots dx_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(e^{t_{i}})e^{\sum_{j=1}^{n}t_{i,j}}dt_{i,1}\cdots dt_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\prod_{i=1}^{N}f_{i}^{{\frac{1}{A\alpha_{i}}}}(e^{t_{i}})e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{\sum_{i=1}^{N}\frac{t_{i,j}}{\alpha_{j}}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
$=\Bigl(\int_{\mathbb{R}^{n}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{As_{j}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}=\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}t^{A}_{j}\Bigr)dt\Bigr)^{A}.$
Remark 6.2.
The statement remains true if we take $(x_{i,1}\cdots x_{i,n})^{\gamma}dx_{i}$, $\gamma>-1,$ for the reference measure instead of the Lebesgue measure.
We prove the main result of this paper.
Theorem 6.3.
(Theorem 1.1) Let $N\geq 2$ be a natural number. Consider the following cost function
$c(x)=c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$
on $(\mathbb{R}^{n})^{N}$. Assume we are given $N$ symmetric convex sets $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ such that $c(x_{1},\cdots,x_{N})\leq 1$ on the set $\prod_{i=1}^{N}A_{i}\subset(\mathbb{R}^{n})^{N}$. Then
$\prod_{i=1}^{N}|A_{i}|\leq|B_{N}|^{N}.$
Let $\{V_{i}\}$, $1\leq i\leq N,$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$. Then
$\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}dy\Bigr)^{N}.$
In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
Proof.
Clearly, $c$-polar transform preserves symmetric and unconditional sets and according to Example 5.2 the cost function satisfies $(j,i_{1},i_{2})$-assumption for all indices. By Corollary 5.4 the desired inequality can be reduced the unconditional case. Thus without loss of generality one can assume that $A_{i}$ are unconditional. Then the set inequality follows from Proposition 6.1 (we take $f_{i}=I_{A_{i}}$ and $\rho(t)=I_{[0,1]}(t)$). The functional inequality follows from Theorem 4.4. The statement about homogeneity of maximizers follows from Corollary 3.4. Finally, description of maximizers in terms of Minkowski functionals follows from Corollary 4.5. ∎
The same arguments can be applied to prove the following result.
Theorem 6.4.
Consider cost functon
$c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}|x_{1,j}|^{\frac{1}{\alpha_{1}}}|x_{2,j}|^{\frac{1}{\alpha_{2}}}\cdots|x_{N,j}|^{\frac{1}{\alpha_{N}}},$
where $\alpha_{i}\leq 1$. Let $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ be $N$ symmetric convex sets satisfying $c(x_{1},\cdots,x_{N})\leq 1$, provided $x_{i}\in A_{i}$, $1\leq i\leq N$. Then
$\prod_{i=1}^{N}|A_{i}|^{\frac{1}{\alpha_{i}}}\leq|B|^{A},$
where $B=\bigl\{t:\sum_{j=1}^{n}|t_{j}|^{A}\leq 1\bigr\}$, $A=\sum_{i=1}^{N}\frac{1}{\alpha_{i}}$.
Let $\{V_{i}\}$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$. Then
$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{A}\sum_{i=1}^{n}|y_{i}|^{A}}dy\Bigr)^{A}.$
In addition, every maximizing tuple $\{\Phi_{i}\}$ with $\Phi_{i}(0)=0$ consists of $A$-homogeneous functions of the form $\{\frac{\frac{1}{\alpha_{i}}}{A}\|x_{i}\|^{A}_{K_{i}}\}$, where $\{K_{i}\}$ is a maximizing tuple of the corresponding set problem.
7. Concluding remarks and open questions
In view of Theorems 1.1, 6.4 it is natural to ask about examples of the Blaschke–Santaló inequality for non-Lebesgue reference measures. One such example can be constructed from the result of Theorem 6.4 just by change of variables
$x_{i,j}={\rm{sign}}(y_{i,j})|y_{i,j}|^{\tau_{i}},\ \ \ {\tau_{i}}\neq 0.$
Examples of maximizers for the functional
(31) $|A|^{\frac{1}{p}}\Bigl(\int_{A^{\circ}}\rho dy\Bigr)^{\frac{1}{q}},$
where $p>1,\frac{1}{p}+\frac{1}{q}=1$ and $\rho$ is $\frac{2-p}{p-1}$-homogeneous can be found in [11]. Another remarkable example is given in Fradelizi–Meyer–Zvavitch [20] (Theorem 7):
(32) $\mu(A)\mu(A^{\circ})\leq\mu(B_{2})^{2}$
for any convex symmetric $A$ and unconditional log-concave measure $\mu$. We remind the reader, that Cordero-Erausquin conjectured in [12] that (32) is true for any symmetric log-concave $\mu$. It is possible to extend the result of [20] to the case of $N>2$ sets and functionals of the type $\prod_{i=1}^{N}\mu(A_{i})$ and the cost function $c=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$ using symmetrization and an adapted variant of Proposition 6.1, but we don’t pursue this here. Unlike this, it is problematic to extend inequality (31) to the case of $N>2$ sets using symmetrization. This is because we need log-concavity of measures for symmetrization and this assumption does not fit well with homogeneity.
Finally, we observe that the barycentric cost function $c_{bar}(x)=\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ does not satisfy the key assumption of $(p_{1},\cdots,p_{N})$-homogeneity and our results can not be applied to this case. We know by the result of Nakamura and Tsuji [32] that the corresponding functional $\prod_{i=1}^{N}\int e^{-V_{i}}dx_{i}$ with constraints $\sum_{i=1}^{N}V_{i}(x_{i})\geq\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ admits $2$-homogeneous (even quadratic) extremizers, but nothing is known about solutions to the corresponding set problem and about the relations between both problems (uniqueness, homogeneity of any maximizing tuple etc).

### v2

@@
 We study Blaschke–Santaló-type inequalities for $N\geq 2$ sets (functions) and a special class of cost functions.
-In particular, we prove new results about reduction of the maximization problem for the Blaschke–Santaló-type functional to homogeneous case (functional inequalities on the sphere) and extend the symmetrization argument to the case of $N>2$ sets.
+In particular, we prove new results about reduction of the maximization problem for the Blaschke–Santaló-type functional to the homogeneous case (functional inequalities on the sphere) and prove new inequalities for $N>2$ sets by symmetrization argument.
 We also discuss links to the multimagrinal optimal transportation problem and the related sharp transportation-information inequalities.
@@
 $x_{i}=(x_{i,1},\cdots,x_{i,j},\cdots,x_{i,n}).$
-We prove the following generalization of the both versions of the Blaschke–Santaló inequality.
+The following interesting example of the Blaschke–Santaló inequality was proved by Kalantzopoulos and Saroglou in [23]
 Theorem 1.1.
-Let $N\geq 2$ be a natural number. Consider the following cost function
+(Kalantzopoulos, Saroglou, [23]) Let $N\geq 2$ be a natural number. Consider the following cost function
 $c(x)=c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$
@@
 $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}dy\Bigr)^{N}.$
-In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
+We extend this result in section 6 of our work, where we get, in particular a generalization for functionals of the type $\prod_{i=1}^{N}|A_{i}|^{\frac{1}{\alpha_{i}}}$ .
 This example can be considered as an illustration of several results of this work, where the central question is the relation between the functional and the set version of the Blaschke–Santaló type functional.
@@
     There are several extensions of the Blaschke–Santaló inequality for special types of the generalized Legendre transform.
-See, for instance, Rotem [35], Artstein–Slomka [3], Artstein–Sadowski–Wyczesany [2], Fradelizi–Gozlan–Sadovsky–Zugmeyer [18].
+See, for instance, Rotem [35], Artstein–Slomka [3], Artstein–Sadovski–Wyczesany [2], Fradelizi–Gozlan–Sadovsky–Zugmeyer [18].
   • Relation to optimal transportation and information theory.
@@
 Some generalizations of this result were obtained by Kalantzopoulos and Saroglou [23].
-In particular, they conjectured the following set version of (6):
+In particular, the following version of (6) was conjectured:
     (7) $\prod_{i=1}^{N}|A_{i}|\leq|B_{2}|^{N},$
@@
 $c(x_{1},\cdots,tx_{i},\cdots,x_{N})=t^{p_{i}}c(x).$
-Another central result of this work is the following Theorem about homogeneity of maximizers of the Blaschke–Santaló functional.
+The central result of this work is the following Theorem about homogeneity of maximizers of the Blaschke–Santaló functional.
 Theorem 1.3.
@@
 In section 2 we discuss the relation of the Blaschke–Santaló inequality to the optimal transportation problem.
+Theorem 1.3 is proved in section 3.
 We show in sections 4-5 that $(p_{1},\cdots,p_{N})$-homogeneous cost functions admit other nice properties.
 In particular, we prove that the corresponding functional and set versions of the Blaschke–Santaló inequality for these cost functions are equivalent.
-In addition, we show that the Blaschke–Santaló functional increases under the appropriate version of the Steiner symmetrization procedure.
+In addition, we discuss the monotonicity of the Blaschke–Santaló functional under the appropriate version of the Steiner symmetrization procedure.
+In section 6 we consider examples.
 Acknowledgements
-The author was supported by RSF project 25-11-00007. The author thanks Matthieu Fradelizi for his interest and comments about the bibliography.
+The author was supported by RSF project 25-11-00007. The author thanks Matthieu Fradelizi for his interest and comments about the bibliography and Christos Saraglou for comments about Theorem 1.1.
 2. Optimal transportation approach
@@
 Let $c\geq 0$ be sufficiently regular cost function on $(\mathbb{S}^{n-1})^{N}$, $\varphi_{i}$ be nonnegative functions satisfying (27) and, in addition, $\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}<\infty$ for all $1\leq i\leq N$.
-Consider mass transportation problem with cost function $\log c$ and probability measures
+Consider the mass transportation problem with the cost function $\log c$ and probability measures
 $\mu_{i}=\frac{1}{\int_{\mathbb{S}^{n-1}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}}\frac{\rho_{i}(\theta_{i})d\theta_{i}}{\varphi^{n+r_{i}}_{i}(\theta_{i})}.$
@@
 4. Inequalities for sets and for functions
-In this section we continue our study of homogeneous solutions and find a direct link between the functional and the set inequality. Let us specify assumptions on $c$ in this section.
+In this section we continue our study of homogeneous solutions and find a direct link between the functional and the set inequality.
+This was partially done in the previous sections, where we find conditions ensuring homogeneity of the maximizers in the functional inequality.
+Here we discuss how to recover the functional inequality from the set inequality using ”lawer-cake” representation.
+Assumption on reference measures:
+$m_{i}=\rho_{i}(x_{i})dx_{i},$
+where $\rho_{i}$ is $r_{i}$-homogeneous.
 Assumption on $c$:
@@
 In this section we consider a generalized symmetrization procedure for $N>2$ functions possessing property to increase the Blaschke–Santaló functional.
+Remark 5.1.
+Unlike sections 3-4 we don’t assume homogeneity of the cost function $c$ and the reference measures.
+Our approach closely follows the arguments of Kalantzopoulos and Saraglou from [23], in particular, our important $(j,i_{1},i_{2})$-assumption below is satisfied by the cost functions considered in [23].
+However we establish our result in a more general situation.
+In particular, we consider more general class of functions and the reference measures.
 Let $c(x_{1},\cdots,x_{N})$ be a continuous function of $N$ variables $x_{i}\in\mathbb{R}^{n}$.
@@
 provided $x_{i}\in A_{i}$ for all $1\leq i\leq N$.
-We will assume in this section that the polar $c$-transform preserves symmetric and convex sets, i.e.
-$B_{j}=\{x_{j}:c(x_{1},\cdots,x_{N})\leq 1,\forall x_{i}\in B_{i},i\neq j\}$
-is symmetric (convex) if $B_{i}$ are symmetric (convex) for all $j\neq i$. Recall that this is true if $c$ satisfies assumption (2) of Corollary 3.4.
 Let us describe the symmetrization with respect to a coordinate unit vector $e_{j}$, $1\leq j\leq n$ for a given couple of indices $i_{1}\neq i_{2}$, $1\leq i_{1},i_{2}\leq N$.
@@
 We say that the cost function $c$ satisfies $(j,i_{1},i_{2})$-assumption if the following holds:
+  1. (1)
+    The polar $c$-transform preserves symmetric and convex sets, i.e.
+    $B_{j}=\{x_{j}:c(x_{1},\cdots,x_{N})\leq 1,\forall x_{i}\in B_{i},i\neq j\}$
+    is symmetric (convex) if $B_{i}$ are symmetric (convex) for all $j\neq i$. Recall that this is true if $c$ satisfies assumption (2) of Corollary 3.4.
+  2. (2)
     The function
-    $(y_{i_{1}},t_{i_{2}})\to c((y_{1},t_{1}),\cdots,(y_{N},t_{N}))$
-    is convex for all fixed $y_{i_{k}}$, $k\neq 1$ and $t_{i_{j}}$, $j\neq 2$
+    $c_{i_{1},i_{2}}(y_{i_{1}},t_{i_{2}})=c((y_{1},t_{1}),\cdots,(y_{N},t_{N}))$
+    has convex level sets for all fixed $y_{i_{k}}$, $k\neq 1$ and $t_{i_{j}}$, $j\neq 2$
+  3. (3)
     $c((y_{1},t_{1}),$ $\cdots,(y_{i_{1}},-t_{i_{1}}),\cdots,(y_{i_{2}},t_{i_{2}}),\cdots(y_{N},t_{N}))$
@@
 for all $(y_{1},t),(y_{1},s)\in A_{1},(y_{i},t_{i})\in A_{i},i>2$. Hence
-$1$ $\geq\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
-$=\frac{1}{2}\Bigl[c\bigl((y_{1},t),(a,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)+c\bigl((y_{1},-s),(b,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr)\Bigr]$
-$\geq c\bigl((y_{1},\frac{t-s}{2}),(\frac{a+b}{2},r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr).$
+$c\bigl((y_{1},-s),(b,r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})=c\bigl((y_{1},s),(b,-r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\leq 1$
+by item (3) of $(n,1,2)$ assumption (see Definition 5.2), hence by item (2)
+$1\geq c\bigl((y_{1},\frac{t-s}{2}),(\frac{a+b}{2},r),(y_{3},t_{3}),\cdots,(y_{N},t_{N})\bigr).$
 Thus $\frac{a+b}{2}\in B(r)$ and the proof is completed. ∎
@@
 6. Main example
+In this section we discuss examples of the Blaschke–Santaló-type inequalities, generalizing Theorem 1.1 from [23].
+Our extensions go in the following directions: 1) extension to unconditional log-concave reference measures in the spirit of Theorem 7 from [20], 2) extension to product functionals with different weights, i.e.
+functionals of the type $\prod_{i=1}^{N}m^{\frac{1}{\alpha_{i}}}(A_{i})$.
+For the case $N=2$ the following result was obtained in Fradelizi–Meyer–Zvavitch [20] (Theorem 7):
+(31) $\mu(A)\mu(A^{\circ})\leq\mu(B_{2})^{2}$
+for any convex symmetric $A$ and unconditional log-concave measure $\mu$. We remind the reader, that Cordero-Erausquin conjectured in [12] that (31) is true for any symmetric log-concave $\mu$.
 Proposition 6.1.
@@
 $\prod_{i=1}^{N}f^{\frac{1}{\alpha_{i}}}_{i}(x_{i})\leq\rho(c(x))$
-for some Borel function $\rho$. Then
-$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(x_{i})dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\bigl(\sum_{j=1}^{n}t^{A}_{j}\bigr)dt\Bigr)^{A},$
+for some Borel function $\rho$. Let $m=e^{-V}dx$ be a measure on $\mathbb{R}^{n}_{+}$ such that the function $(t_{1},\cdots,t_{n})\to V(e^{t_{1}},\cdots,e^{t_{n}})$ is convex.
+Then
+$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(x_{i})m(dx_{i})\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\bigl(\sum_{j=1}^{n}t^{A}_{j}\bigr)m(dt)\Bigr)^{A},$
 where $A=\sum_{j=1}^{N}\frac{1}{\alpha_{j}}$.
 In particular, if functions $V_{i}(x_{i})$ satisfy $\sum_{i=1}^{N}V_{i}(x_{i})\geq\sum_{j=1}^{n}x^{\frac{1}{\alpha_{1}}}_{1,j}\cdots x^{\frac{1}{\alpha_{N}}}_{N,j}$, then
-$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}e^{-\frac{1}{A}\sum_{j=1}^{n}t^{A}_{j}}dt\Bigr)^{A}.$
+$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}e^{-\alpha_{i}V_{i}(x_{i})}m(dx_{i})\Bigr]^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}_{+}}e^{-\frac{1}{A}\sum_{j=1}^{n}t^{A}_{j}}m(dt)\Bigr)^{A}.$
 Proof.
 Following [19] we apply the exponential change of variables : $x_{i,j}=e^{t_{i,j}}$ (or simply $x_{i}=e^{t_{i}}$) and the Prekopa–Leindler inequality
-$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})dx_{i,1}\cdots dx_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(e^{t_{i}})e^{\sum_{j=1}^{n}t_{i,j}}dt_{i,1}\cdots dt_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
-$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\prod_{i=1}^{N}f_{i}^{{\frac{1}{A\alpha_{i}}}}(e^{t_{i}})e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
-$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{\sum_{i=1}^{N}\frac{t_{i,j}}{\alpha_{j}}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
-$=\Bigl(\int_{\mathbb{R}^{n}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{As_{j}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}=\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}t^{A}_{j}\Bigr)dt\Bigr)^{A}.$
-Remark 6.2.
-The statement remains true if we take $(x_{i,1}\cdots x_{i,n})^{\gamma}dx_{i}$, $\gamma>-1,$ for the reference measure instead of the Lebesgue measure.
-We prove the main result of this paper.
+$\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})e^{-V(x_{i})}dx_{i}\Bigr]^{\frac{1}{\alpha_{i}}}$ $=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}}f_{i}(x_{i})e^{-V(x_{i})}dx_{i,1}\cdots dx_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
+$=\prod_{i=1}^{N}\Bigl[\int_{\mathbb{R}^{n}_{+}}f_{i}(e^{t_{i}})e^{-V(e^{t_{i}})}e^{\sum_{j=1}^{n}t_{i,j}}dt_{i,1}\cdots dt_{i,n}\Bigr]^{\frac{1}{\alpha_{i}}}$
+$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\prod_{i=1}^{N}f_{i}^{{\frac{1}{A\alpha_{i}}}}(e^{t_{i}})e^{-\frac{1}{A\alpha_{i}}V(e^{t_{i}})}e^{\sum_{k=1}^{n}s_{k}}ds_{1}\cdots ds_{n}\Bigr)^{A}$
+$\leq\Bigl(\int_{\mathbb{R}^{n}}\sup_{s_{k}=\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i,k}}{\alpha_{i}}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{\sum_{i=1}^{N}\frac{t_{i,j}}{\alpha_{j}}}\Bigr)e^{-V(e^{\frac{1}{A}\sum_{i=1}^{N}\frac{t_{i}}{\alpha_{i}}})}e^{\sum_{k=1}^{n}s_{k}}ds\Bigr)^{A}$
+$=\Bigl(\int_{\mathbb{R}^{n}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}e^{As_{j}}\Bigr)e^{\sum_{k=1}^{n}s_{k}}e^{-V(e^{s})}ds\Bigr)^{A}=\Bigl(\int_{\mathbb{R}^{n}_{+}}\rho^{\frac{1}{A}}\Bigl(\sum_{j=1}^{n}t^{A}_{j}\Bigr)e^{-V(t)}dt\Bigr)^{A}.$
 Theorem 6.2.
-(Theorem 1.1) Let $N\geq 2$ be a natural number. Consider the following cost function
+(Extension of Theorem 1.1) Let $N\geq 2$ be a natural number and $m$ be a log-concave unconditional measure on $\mathbb{R}^{n}$. Consider the following cost function
 $c(x)=c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$
@@
 Then
-$\prod_{i=1}^{N}|A_{i}|\leq|B_{N}|^{N}.$
-Let $\{V_{i}\}$, $1\leq i\leq N,$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$.
+$\prod_{i=1}^{N}m(A_{i})\leq m(B_{N})^{N}.$
+Let $\{V_{i}\}$, $1\leq i\leq N,$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(m)$ for every $i$.
 Then
-$\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}dy\Bigr)^{N}.$
-In addition, let $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
+$\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}m(dx_{i})\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{N}\sum_{i=1}^{n}|y_{i}|^{N}}m(dy)\Bigr)^{N}.$
+In addition, let $m$ be the Lebesgue measure and $\{\Phi_{i}\}$ be a maximizing tuple of the functional $\prod_{i=1}^{N}\int_{\mathbb{R}^{n}}e^{-V_{i}(x_{i})}dx_{i}$ under the same constraints with $\Phi_{i}(0)=0$.Then $\Phi_{i}=\frac{1}{N}\|x_{i}\|^{N}_{K_{i}}$, where $\{K_{i}\}$ is a maximizing tuple for the corresponding set problem and $\|\cdot\|_{K_{i}}$ are the associated Minkowski functionals.
 Proof.
 Clearly, $c$-polar transform preserves symmetric and unconditional sets and according to Example 5.3 the cost function satisfies $(j,i_{1},i_{2})$-assumption for all indices.
-By Corollary 5.4 the desired inequality can be reduced the unconditional case.
+By Corollary 5.5 the desired inequality can be reduced to the unconditional case.
 Thus without loss of generality one can assume that $A_{i}$ are unconditional.
-Then the set inequality follows from Proposition 6.1 (we take $f_{i}=I_{A_{i}}$ and $\rho(t)=I_{[0,1]}(t)$).
-The functional inequality follows from Theorem 4.4.
+The same symmetrization procedure can be done for functions (we don’t present the complete proof here, but we can refer to [11] Theorem 4.1, where the symmetrization for functions was considered in detail for $N=2$.
+The case $N>2$ follows the arguments of Lemma 5.4 and Theorem 4.1 of [11]).
+Finally, after reduction to the unconditional case, both inequalities (for sets and functions) follow from Proposition 6.1.
 The statement about homogeneity of maximizers follows from Corollary 3.4.
@@
 $c(x_{1},\cdots,x_{N})=\sum_{j=1}^{n}|x_{1,j}|^{\frac{1}{\alpha_{1}}}|x_{2,j}|^{\frac{1}{\alpha_{2}}}\cdots|x_{N,j}|^{\frac{1}{\alpha_{N}}},$
-where $\alpha_{i}\leq 1$. Let $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ be $N$ symmetric convex sets satisfying $c(x_{1},\cdots,x_{N})\leq 1$, provided $x_{i}\in A_{i}$, $1\leq i\leq N$. Then
-$\prod_{i=1}^{N}|A_{i}|^{\frac{1}{\alpha_{i}}}\leq|B|^{A},$
+where $\alpha_{i}\leq 1$.
+Let $A_{i}\subset\mathbb{R}^{n}$, $1\leq i\leq N$ be $N$ symmetric convex sets satisfying $c(x_{1},\cdots,x_{N})\leq 1$, provided $x_{i}\in A_{i}$, $1\leq i\leq N$ and $m$ be a log-concave unconditional measure.
+Then
+$\prod_{i=1}^{N}m(A_{i})^{\frac{1}{\alpha_{i}}}\leq m(B)^{A},$
 where $B=\bigl\{t:\sum_{j=1}^{n}|t_{j}|^{A}\leq 1\bigr\}$, $A=\sum_{i=1}^{N}\frac{1}{\alpha_{i}}$.
-Let $\{V_{i}\}$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(dx_{i})$ for every $i$.
+Let $\{V_{i}\}$ be even measurable functions with values in $(-\infty,+\infty]$ and satisfying $\sum_{i=1}^{N}V_{i}(x_{i})\geq c(x).$ Assume, in addition, that $e^{-V_{i}}\in L^{1}(m)$ for every $i$.
 Then
-$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}dx_{i}\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{A}\sum_{i=1}^{n}|y_{i}|^{A}}dy\Bigr)^{A}.$
-In addition, every maximizing tuple $\{\Phi_{i}\}$ with $\Phi_{i}(0)=0$ consists of $A$-homogeneous functions of the form $\{\frac{\frac{1}{\alpha_{i}}}{A}\|x_{i}\|^{A}_{K_{i}}\}$, where $\{K_{i}\}$ is a maximizing tuple of the corresponding set problem.
-7. Concluding remarks and open questions
-In view of Theorems 1.1, 6.4 it is natural to ask about examples of the Blaschke–Santaló inequality for non-Lebesgue reference measures.
-One such example can be constructed from the result of Theorem 6.4 just by change of variables
+$\prod_{i=1}^{N}\Bigl(\int_{\mathbb{R}^{n}}e^{-\alpha_{i}V_{i}(x_{i})}m(dx_{i})\Bigr)^{\frac{1}{\alpha_{i}}}\leq\Bigl(\int_{\mathbb{R}^{n}}e^{-\frac{1}{A}\sum_{i=1}^{n}|y_{i}|^{A}}m(dy)\Bigr)^{A}.$
+In addition, let $m$ be the Lebesgue measure.
+Then every maximizing tuple $\{\Phi_{i}\}$ with $\Phi_{i}(0)=0$ consists of $A$-homogeneous functions of the form $\{{\frac{1}{A\alpha_{i}}}\|x_{i}\|^{A}_{K_{i}}\}$, where $\{K_{i}\}$ is a maximizing tuple of the corresponding set problem.
+Remark 6.4.
+In view of Theorems 1.1, 6.3 it is natural to ask about examples of the Blaschke–Santaló inequality for homogeneous non-Lebesgue reference measures.
+One such example can be constructed from the result of Theorem 6.3 with Lebesgue reference measure just by change of variables
 $x_{i,j}={\rm{sign}}(y_{i,j})|y_{i,j}|^{\tau_{i}},\ \ \ {\tau_{i}}\neq 0.$
@@
 (32) $|A|^{\frac{1}{p}}\Bigl(\int_{A^{\circ}}\rho dy\Bigr)^{\frac{1}{q}},$
-where $p>1,\frac{1}{p}+\frac{1}{q}=1$ and $\rho$ is $\frac{2-p}{p-1}$-homogeneous can be found in [11]. Another remarkable example is given in Fradelizi–Meyer–Zvavitch [20] (Theorem 7):
-(32) $\mu(A)\mu(A^{\circ})\leq\mu(B_{2})^{2}$
-for any convex symmetric $A$ and unconditional log-concave measure $\mu$.
-We remind the reader, that Cordero-Erausquin conjectured in [12] that (32) is true for any symmetric log-concave $\mu$.
-It is possible to extend the result of [20] to the case of $N>2$ sets and functionals of the type $\prod_{i=1}^{N}\mu(A_{i})$ and the cost function $c=\sum_{j=1}^{n}x_{1,j}x_{2,j}\cdots x_{N,j}$ using symmetrization and an adapted variant of Proposition 6.1, but we don’t pursue this here.
-Unlike this, it is problematic to extend inequality (31) to the case of $N>2$ sets using symmetrization.
+where $p>1,\frac{1}{p}+\frac{1}{q}=1$ and $\rho$ is $\frac{2-p}{p-1}$-homogeneous can be found in [11].
+We observe that it is problematic to extend inequality (32) to the case of $N>2$ sets using symmetrization in the homogeneous case.
 This is because we need log-concavity of measures for symmetrization and this assumption does not fit well with homogeneity.
-Finally, we observe that the barycentric cost function $c_{bar}(x)=\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ does not satisfy the key assumption of $(p_{1},\cdots,p_{N})$-homogeneity and our results can not be applied to this case.
-We know by the result of Nakamura and Tsuji [32] that the corresponding functional $\prod_{i=1}^{N}\int e^{-V_{i}}dx_{i}$ with constraints $\sum_{i=1}^{N}V_{i}(x_{i})\geq\sum_{i\neq j}\langle x_{i},x_{j}\rangle$ admits $2$-homogeneous (even quadratic) extremizers, but nothing is known about solutions to the corresponding set problem and about the relations between both problems (uniqueness, homogeneity of any maximizing tuple etc).